Jefferson Lab HALL A Logbook for April 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 320292 04/30/10 20:23 dalton IT changed for 120 Hz
  • 320291 04/30/10 20:17 Jack (Techs) Pillage first THEN burn (Right Dipole is responding)
  • 320284 04/30/10 19:05 dalton Reversal rate changed to 120 Hz with long T_settle
  • 320280 04/30/10 18:30 M Friend Compton photon detector lead hut
  • 320270 04/30/10 17:42 Jack (Techs) Right Dipole status
  • 320248 04/30/10 15:04 O.Glamazdin Moller magnet has been filled
  • 320244 04/30/10 12:36 paschke gas system alarm / Tech
  • 320234 04/30/10 11:58 R. Michaels raster timing check
  • 320225 04/30/10 11:49 R. Michaels story of cavities (fixed now, we think)
  • 320203 04/30/10 08:38 kelleher Target Moved to Empty Position
  • Thursday

  • 320187 04/29/10 17:29 kelleher NO Running in Hall A tonight
  • 320186 04/29/10 16:59 Abdurahim Compton cavity status
  • 320185 04/29/10 13:21 ZAFAR AHMED Current cavities are satuarting
  • 320184 04/29/10 11:31 ZAFAR AHMED April 29 Day Shift Summary
  • 320183 04/29/10 11:21 Abdurahim Compton cavity lock came back...
  • 320182 04/29/10 11:19 kelleher Shifts Cancelled Until May 1 Swing
  • 320181 04/29/10 11:07 ZAFAR AHMED Restricted access for hall A
  • 320179 04/29/10 10:48 A. Camsonne Compton electron detector hit pattern
  • 320177 04/29/10 10:35 A. Camsonne Compton electron detector threshold scan
  • 320175 04/29/10 10:31 ZAFAR AHMED Current cavity coorelation
  • 320166 04/29/10 09:42 O.Glamazdin Moller magnet has been refilled
  • 320163 04/29/10 09:25 ZAFAR AHMED Controlled access
  • 320162 04/29/10 09:22 A. Camsonne Compton detector threshold check - looks good
  • 320159 04/29/10 08:22 M Friend Compton cavity locking in wrong mode
  • 320158 04/29/10 08:09 V. Sulkosky Owl Shift Summary
  • 320156 04/29/10 08:00 deng stripcharts of target
  • 320155 04/29/10 07:57 D.Wang IHWP put IN
  • 320153 04/29/10 07:54 apar Halfwave plate IN, Slug 100
  • 320151 04/29/10 07:52 M Friend Compton HV back on
  • 320150 04/29/10 07:38 V. Sulkosky Temperatures
  • 320149 04/29/10 07:24 V. Sulkosky No Trips 3 Hours 15'
  • 320144 04/29/10 06:02 Ole Hansen adaql5 often runs out of memory
  • 320143 04/29/10 05:38 V. Sulkosky Raster Test Postponed
  • 320140 04/29/10 05:10 V. Sulkosky Hall A FastFeedback Problems Summary from ELOG
  • 320139 04/29/10 05:04 V. Sulkosky BTA Claiming no ABU
  • 320138 04/29/10 05:02 D.Wang Current and Charge from prompt analysis?
  • 320137 04/29/10 04:37 V. Sulkosky Rad/beam position Charts
  • 320134 04/29/10 04:19 V. Sulkosky Target Vacuum
  • 320133 04/29/10 04:14 V. Sulkosky Lumi HV's Still Off
  • 320132 04/29/10 04:11 D.Wang dithering plots
  • 320131 04/29/10 04:07 V. Sulkosky Hall A Tools Screen
  • 320129 04/29/10 03:34 deng target moved to thick carbon for production
  • 320128 04/29/10 03:33 D.Wang spot++
  • 320126 04/29/10 03:30 D.Wang T5 rate 5nA with lead target
  • 320122 04/29/10 03:24 D.Wang T5 rate at 5nA
  • 320121 04/29/10 03:23 Deng Target moved to Lead/Diamond
  • 320120 04/29/10 03:15 Deng Target moved to BeO
  • 320119 04/29/10 03:11 D.Wang spot++
  • 320115 04/29/10 03:00 V. Sulkosky Beam Permit
  • 320114 04/29/10 02:56 V. Sulkosky Beam Maybe soon!
  • 320104 04/29/10 01:31 D.Wang rebooted CODA
  • 320099 04/29/10 00:49 V. Sulkosky Beam Update
  • 320098 04/29/10 00:20 V. Sulkosky Hall Tools
  • 320096 04/29/10 00:01 A. Deur Swing Shift Summary
  • Wednesday

  • 320094 04/28/10 23:57 paschke retry log entry: oversampling and "pinhole" scan in integration window
  • 320093 04/28/10 23:57 paschke retry log entry: oversampling and "pinhole" scan in integration window
  • 320092 04/28/10 23:49 adaq 2nd Vacuum jump
  • 320091 04/28/10 23:29 Carlos Target snashot
  • 320090 04/28/10 23:25 R. Michaels raster unsynched, then no beam.
  • 320089 04/28/10 22:54 A. Camsonne Plot for the Compton electron spectrum are in entry 319934
  • 320088 04/28/10 22:53 mercado Turned off LUMI voltages
  • 320087 04/28/10 22:47 A. Camsonne GSO HV tripped
  • 320086 04/28/10 22:46 mercado Check correlations at 50uA on PREX target
  • 320083 04/28/10 22:30 A. Camsonne Compton GSO on at 10 uA and 50 uA
  • 320082 04/28/10 22:28 R. Michaels Pb noise tests, raster synched.
  • 320079 04/28/10 22:17 mercado Checking correlations at 10uA on PREX target
  • 320077 04/28/10 22:06 ZAFAR AHMED Cavity Coorelation
  • 320070 04/28/10 21:43 Carlos Huge vacuum jump when beam on lead
  • 320067 04/28/10 21:37 A. Deur Laser attenuator/Slit values for ~5nA
  • 320066 04/28/10 21:35 R. Michaels spot++ for Pb/D. Hole
  • 320060 04/28/10 21:08 A. Deur Raster size
  • 320058 04/28/10 20:55 ZAFAR AHMED Cavity Pregains and Phases for 50 uA.
  • 320057 04/28/10 20:35 A. Deur Septum on
  • 320052 04/28/10 19:00 O.Glamazdin Moller results
  • 320051 04/28/10 18:19 dalton Slugs moved
  • 320050 04/28/10 18:08 Abdurahim Compton cavity power history plot
  • 320049 04/28/10 17:58 jwexler Changed feedback curmon device to bcm3
  • 320039 04/28/10 17:42 M Friend Compton HV off
  • 320036 04/28/10 17:36 M Friend Compton FADC DAQ to 240Hz
  • 320035 04/28/10 17:35 R.Michaels re: 15V supply on cavities (no cavities since Sunday afternoon)
  • 320033 04/28/10 17:26 paschke green monster doesn't update display automatically: timeboard parameters may display incorrect after coda download
  • 320030 04/28/10 17:23 M Friend EDET to garage
  • 320026 04/28/10 17:16 M Friend EDET to -3mm
  • 320025 04/28/10 16:24 ZAFAR AHMED Cavity oscillator power supply is on
  • 320024 04/28/10 16:09 M Friend EDET to garage for access
  • 320023 04/28/10 15:54 M Friend EDET to -3 mm
  • 320022 04/28/10 14:29 nanda Electron Spectrum
  • 320021 04/28/10 14:16 M Friend Compton FADC DAQ to 30Hz
  • 320020 04/28/10 14:08 R. Michaels re: timing board parameters
  • 320019 04/28/10 13:46 Abdurahim dithering related problem
  • 320016 04/28/10 13:38 mercado Cavity BCMs still not working
  • 320015 04/28/10 13:31 mercado IT was not set properly for Injector
  • 320014 04/28/10 13:26 mercado Feedback did not work this morning
  • 320013 04/28/10 13:18 Abdurahim bcmcav2 & bcmcav3 ASym
  • 320012 04/28/10 13:09 mercado Dither related glitch
  • 320011 04/28/10 12:11 Abdurahim a strange A_t detector histograms
  • 320010 04/28/10 11:33 O.Glamazdin BPM 1HO1 reading is very noisy
  • 320009 04/28/10 10:21 O.Glamazdin Helicity board changes for Moller
  • 320008 04/28/10 10:21 Fansler/folt target vacuum
  • 320005 04/28/10 09:52 O.Glamazdin Moller magnet has been refilled
  • 320004 04/28/10 09:47 Ole Hansen adaql8 display/mouse fixed
  • 320002 04/28/10 09:47 R.M., W. Gun raster synched to helicity flip
  • 319997 04/28/10 09:33 ZAFAR AHMED Target is empty target.
  • 319996 04/28/10 09:29 romanip Rebooting Moller VME Crate
  • 319993 04/28/10 08:49 R. Michaels re: how to setup and check cavities
  • 319990 04/28/10 08:16 W. Deconinck adaql8: no graphical interface, mouse problem
  • 319989 04/28/10 08:16 Ole Hansen adaql8 mouse
  • 319987 04/28/10 08:02 W. Deconinck Rebooting adaql8
  • 319986 04/28/10 07:57 jwexler Shift summary - Owl 4/28/2010
  • 319983 04/28/10 07:37 jwexler Charge feedback problems
  • 319978 04/28/10 05:32 jwexler LATE ENTRY - MCC has difficulties with test.
  • 319977 04/28/10 05:16 jwexler Hall A tools GUI.
  • 319976 04/28/10 05:15 jwexler Radiation and temperature stripcharts
  • 319975 04/28/10 05:14 jwexler Target GUI and stripchart
  • 319971 04/28/10 04:47 jwexler Changed slug number - Slug 98
  • 319970 04/28/10 04:46 apar Halfwave plate OUT, Slug 98
  • 319969 04/28/10 04:42 jwexler Target changed to thick carbon
  • 319968 04/28/10 04:32 jwexler MCC starting their test.
  • 319967 04/28/10 04:31 jwexler Target to Empty
  • 319966 04/28/10 04:23 jwexler Changed bcm/det pedestals in db file.
  • 319962 04/28/10 03:51 jwexler Inj. IT improperly set
  • 319957 04/28/10 03:35 jwexler Changed to IT = 1535
  • 319944 04/28/10 00:20 ZAFAR AHMED Online GUI for Cavity Monitors
  • 319943 04/28/10 00:16 pking The fast excursions in INJ data also occur when beam is off.
  • 319942 04/28/10 00:11 jwexler Target IOC rebooted
  • Tuesday

  • 319941 04/27/10 23:59 M Friend Compton EtrocOnly runs
  • 319940 04/27/10 23:47 Carlos Swing shift summary
  • 319939 04/27/10 23:45 A. Camsonne Compton electron detector in garage position
  • 319934 04/27/10 23:42 A. Camsonne Electron detector Compton spectrum
  • 319932 04/27/10 23:38 Carlos Vacuum evolution
  • 319931 04/27/10 23:26 pking Fast spikes observed in INJ data
  • 319930 04/27/10 22:57 Carlos IHWP out
  • 319929 04/27/10 22:52 jhlee *** IHWP OUT ***
  • 319927 04/27/10 22:48 apar Halfwave plate OUT, Slug 97
  • 319925 04/27/10 22:37 A. Camsonne Compton electron detector in position 3 mm
  • 319924 04/27/10 22:34 ZAFAR AHMED Cahnge of gain of current cavities
  • 319923 04/27/10 22:28 ZAFAR AHMED Current cavities gain changed during run 3241
  • 319922 04/27/10 22:23 mcanan Target Screen Shots
  • 319921 04/27/10 22:10 A. Camsonne Requested to mask Compton electron detector FSD to reboot front end
  • 319918 04/27/10 21:51 M Friend Compton FADC thresh changed
  • 319915 04/27/10 21:42 M Friend Comptno FADC crl accum thresh changed
  • 319912 04/27/10 21:27 M Friend Compton asymmetry sign flipped
  • 319909 04/27/10 20:34 M Friend Compton asymmetry
  • 319907 04/27/10 19:51 jhlee lumi screenshots after changing HVs
  • 319905 04/27/10 19:48 dalton Changed 2 lumi HV
  • 319904 04/27/10 19:25 mcanan vacuum pressure of the scattering chamber
  • 319903 04/27/10 19:19 jhlee screenshot for lumis
  • 319902 04/27/10 19:18 jhlee dithering screenshot
  • 319901 04/27/10 19:15 Carlos Chamber vacuum worse again
  • 319899 04/27/10 18:57 mcanan Target Motion
  • 319898 04/27/10 18:52 M Friend Compton vertical scan
  • 319897 04/27/10 18:52 jhlee spot++
  • 319893 04/27/10 18:43 M Friend Compton rates
  • 319891 04/27/10 17:59 dalton HWP is still IN, we performed a WIEN flip
  • 319883 04/27/10 17:24 apar Halfwave plate IN, Slug 96
  • 319879 04/27/10 16:32 A. Camsonne GEM HV off
  • 319878 04/27/10 16:03 Saha/jpchen shift summary
  • 319876 04/27/10 15:55 jpchen target vacuum and temperature charts
  • 319875 04/27/10 15:53 jpchen dithering off
  • 319873 04/27/10 15:02 jpchen requested beam permit in hall
  • 319872 04/27/10 15:00 jpchen Target to EMPTY
  • 319870 04/27/10 15:00 jpchen target ioc rebooted
  • 319867 04/27/10 13:51 mercado Sync Monitor Officially Working
  • 319866 04/27/10 13:46 Kiadtisak GEM's HV revised
  • 319863 04/27/10 13:34 A. Camsonne GEM test done- source out - Left HRS closed
  • 319840 04/27/10 12:00 Abdurahim BCM cav Asymmetry
  • 319836 04/27/10 11:45 Abdurahim Right A_T detector abnormality
  • 319835 04/27/10 11:43 folts target vacuum/tech
  • 319834 04/27/10 11:43 A. Camsonne GEM HV changed
  • 319832 04/27/10 11:40 jpchen temerature and radiation charts
  • 319830 04/27/10 11:34 jpchen Controlled access activities
  • 319828 04/27/10 11:31 A. Camsonne Gem HV on
  • 319827 04/27/10 11:29 O.Glamazdin Moller magnet has been filled
  • 319826 04/27/10 11:26 A. Camsonne GEM source test - No one allowed in the Left HRS without radcon approval
  • 319824 04/27/10 10:43 jpchen Septum off for controlled access
  • 319823 04/27/10 10:10 jpchen scattering chamber vacuum stable at 1.5*10-5
  • 319822 04/27/10 10:00 silwal Oversample data from the qweak adcs on the injector bcm
  • 319821 04/27/10 09:49 jpchen scattering chamber vacuum alarm
  • 319813 04/27/10 09:41 ZAFAR AHMED Cavities gain min and max test(2)
  • 319801 04/27/10 08:50 ZAFAR AHMED Cavities at min and max gains.
  • 319799 04/27/10 08:01 jwexler Shift summary, 04/27/10
  • 319791 04/27/10 07:49 jwexler Compton HV trip
  • 319789 04/27/10 07:38 jwexler Prompt & PANGUIN misreading beam current
  • 319784 04/27/10 05:59 jwexler Compton HV trip
  • 319782 04/27/10 05:23 jwexler ROC23 reset.
  • 319778 04/27/10 04:37 jwexler Rad. and temp. strip charts
  • 319777 04/27/10 04:35 jwexler Target GUI and stripchart
  • 319776 04/27/10 04:32 jwexler Hall A tools
  • 319775 04/27/10 04:29 jwexler Detectors and lumis saturated.
  • 319768 04/27/10 03:30 jwexler Asymmetries for different ramp delay settings.
  • 319767 04/27/10 03:01 jwexler Changed bcm and det. pedestals in db file.
  • 319754 04/27/10 01:58 paschke RHWP 2900 for large charge asymmetry, run 3210
  • 319752 04/27/10 01:54 dalton Oversample charge asymmetry measurement
  • 319737 04/27/10 01:07 jwexler Compton HV trip
  • 319731 04/27/10 00:26 mercado Pedestal change to database for cavities
  • 319730 04/27/10 00:03 ZAFAR AHMED Cavities signal is good in RHRS but not in Parity DAQ.
  • 319728 04/27/10 00:00 A. Deur 04/26/2010 Swing Shift summary
  • Monday

  • 319708 04/26/10 22:39 Abdurahim Compton S/B ratio is 0.07
  • 319701 04/26/10 22:16 benesch beam should likely be higher or tilted down in Compton
  • 319696 04/26/10 21:43 paschke createDB not always picking up on oversample setting
  • 319686 04/26/10 21:02 benesch S/N = 0.07 for Compton
  • 319679 04/26/10 20:47 jhlee oversampling number is wrong
  • 319672 04/26/10 20:21 R. Michaels re: bad cavities
  • 319668 04/26/10 19:37 jhlee prompt dosen't go
  • 319667 04/26/10 19:35 Abdurahim Compton Cavity has been restored
  • 319662 04/26/10 18:09 jhlee bcmcav3 strange signal
  • 319654 04/26/10 17:28 jhlee CODA crashed, restarted it
  • 319652 04/26/10 17:23 jpchen Target alarms should be ok now
  • 319650 04/26/10 17:13 jhlee lumi 3
  • 319649 04/26/10 17:11 ITARD Target Screens
  • 319648 04/26/10 16:50 A. Deur Hall A tools
  • 319647 04/26/10 16:45 mercado Bad Cavity Signals Again
  • 319646 04/26/10 16:42 jhlee Dithering Screenshot
  • 319644 04/26/10 16:18 O.Glamazdin Moller magnet has been refilled
  • 319642 04/26/10 15:44 W. Deconinck Beam off for Moller refill
  • 319640 04/26/10 15:32 W. Deconinck Called MCC for escorted access
  • 319639 04/26/10 15:19 W. Deconinck Target tools screen capture
  • 319638 04/26/10 15:15 W. Deconinck Compton photon detector trip
  • 319637 04/26/10 15:00 Abdurahim Double peak in BPM4bx
  • 319636 04/26/10 14:53 W. Deconinck Compton photon detector trip
  • 319635 04/26/10 14:49 hisham IHWP is IN
  • 319634 04/26/10 14:49 Abdurahim BCM cav Asym plots ?
  • 319633 04/26/10 14:48 W. Deconinck Good production run list for thick carbon
  • 319631 04/26/10 14:43 apar Halfwave plate IN, Slug 95
  • 319627 04/26/10 14:16 Abdurahim Srange BCM cav asymmetry
  • 319626 04/26/10 13:49 dalton Reprompting slug 13
  • 319625 04/26/10 13:41 dalton Changes to BCM datamap
  • 319622 04/26/10 12:59 mercado Added Injector BCM and changed raster pickoff currents
  • 319621 04/26/10 12:59 Abdurahim Strange looking BCM cavity asymmetry
  • 319620 04/26/10 12:49 mercado Full BCM Datamap
  • 319617 04/26/10 12:33 mercado Full BCM Datamap
  • 319616 04/26/10 12:07 jpchen target false alarm
  • 319615 04/26/10 11:59 Abdurahim Strange BPM asymmetry
  • 319614 04/26/10 11:57 Abdurahim Large BCM asymmetry, bcm cavity monitor calibration error ??
  • 319611 04/26/10 11:22 M Friend Compton EPICS readback changed
  • 319608 04/26/10 10:40 W. Deconinck Photon detector HV trip
  • 319605 04/26/10 09:54 riordan autosnaps restarted
  • 319597 04/26/10 09:15 W. Deconinck Compton DAQ started again after cavity fixed
  • 319594 04/26/10 09:02 W. Deconinck Runbird now warns if not on adaql5
  • 319593 04/26/10 08:47 W. Deconinck runbird should be started on adaql5
  • 319592 04/26/10 08:40 hisham Feedback monitor
  • 319589 04/26/10 08:16 hisham Dithering
  • 319588 04/26/10 08:13 Mindy Hall-A general tool
  • 319585 04/26/10 08:07 paschke feedback on bcm1, curmon on bcm3... changed low beam, but need to normalize this
  • 319584 04/26/10 08:06 jwexler Shift Summary - Owl 04/26/2010
  • 319583 04/26/10 08:05 W. Deconinck Compton laser cavity lock lost
  • 319581 04/26/10 07:56 paschke charge feedback not working
  • 319577 04/26/10 06:48 jwexler IHWP now OUT, slug 16.
  • 319576 04/26/10 06:46 apar Halfwave plate OUT, Slug 16
  • 319572 04/26/10 06:05 jwexler Compton trip
  • 319571 04/26/10 05:20 jwexler Compton trip
  • 319568 04/26/10 05:02 jwexler Hall A Tools screengrab
  • 319567 04/26/10 05:02 jwexler Radiation and temperature stripcharts
  • 319566 04/26/10 05:01 jwexler Target GUI and stripchart
  • 319561 04/26/10 03:10 jwexler Possible dithering-panguin issue
  • 319560 04/26/10 02:47 jwexler Dithering issue
  • 319557 04/26/10 02:35 jwexler Hall A tools
  • 319556 04/26/10 02:26 jwexler Target GUI and strip chart
  • 319555 04/26/10 02:22 jwexler Radiation and temperature stripcharts
  • 319554 04/26/10 02:21 jwexler No longer in large charge asymmetry settings
  • 319551 04/26/10 02:18 jwexler BCM pedestal calibration from large charge asymmetry runs.
  • 319546 04/26/10 01:54 Apar test
  • 319543 04/26/10 01:31 mercado Zero beam pedestals for 120Hz
  • 319541 04/26/10 01:30 jwexler Large charge asymmetry run
  • 319540 04/26/10 01:28 jwexler Changed rotating waveplate
  • 319538 04/26/10 01:25 Apar Change pedestals for 240Hz
  • 319535 04/26/10 01:16 silwal prompt working fine now
  • 319530 04/26/10 01:00 mercado BCM added to datamap
  • 319528 04/26/10 00:50 R. Michaels adjust blumi3 HV.
  • 319525 04/26/10 00:40 R. Michaels further mods to DAQ for C12
  • 319522 04/26/10 00:32 jwexler Compton HV trip
  • 319521 04/26/10 00:25 jwexler Slug number incremented to 15 for thick carbon production
  • 319520 04/26/10 00:21 apar Halfwave plate IN, Slug 15
  • 319517 04/26/10 00:13 R. Michaels Switch to thick C12
  • 319515 04/26/10 00:07 jwexler Target alarm screen
  • 319514 04/26/10 00:03 LeRose Shift Summary
  • 319513 04/26/10 00:02 jwexler False gas panel target alarm returned.
  • Sunday

  • 319511 04/25/10 23:59 jwexler Target IOC crash.
  • 319509 04/25/10 23:54 jwexler Target changed - thick carbon
  • 319503 04/25/10 22:29 jhlee Target Strip Chart
  • 319502 04/25/10 22:24 Abdurahim Compton
  • 319501 04/25/10 22:11 M Friend Compton crl thresholds changed
  • 319498 04/25/10 21:22 LeRose Radmon & Temps
  • 319495 04/25/10 21:00 jhlee prompt *** Break *** segmentation violation
  • 319494 04/25/10 20:36 Jack VDC HV control
  • 319493 04/25/10 20:26 jhlee dithering screenshot of PANGUIN
  • 319490 04/25/10 19:58 ZAFAR AHMED Target Chart
  • 319489 04/25/10 19:52 R. Michaels re: recovering VDCs (late entry)
  • 319488 04/25/10 19:25 LeRose Temp & Radmon
  • 319486 04/25/10 19:21 apar Halfwave plate IN, Slug 14
  • 319484 04/25/10 19:17 ZAFAR AHMED Vacuum Pressure No alaram
  • 319481 04/25/10 18:56 mercado DAQ sync signals at 240Hz
  • 319480 04/25/10 18:55 ZAFAR AHMED Beam Recovery Raster at 5 uA. Lead Diamond 3
  • 319474 04/25/10 18:45 ZAFAR AHMED Beam Recovery Spot size Lead Diamond
  • 319471 04/25/10 18:40 ZAFAR AHMED Beam Recovery - Lead Diamond # 3 Rates
  • 319470 04/25/10 18:36 ZAFAR AHMED Target is Lead Diamon # 3
  • 319469 04/25/10 18:34 ZAFAR AHMED Slit is 12.7 and Attn. is 65
  • 319468 04/25/10 18:33 ZAFAR AHMED Beam Recovery - Left Upper Scint Rates at 5 nA
  • 319467 04/25/10 18:21 ZAFAR AHMED Beam Recovery - Target is BeO
  • 319466 04/25/10 18:16 ZAFAR AHMED Beam Recovery - Spot size empty target
  • 319463 04/25/10 18:13 ZAFAR AHMED Beam current and position
  • 319457 04/25/10 17:59 LeRose Septum Magnet Restored
  • 319455 04/25/10 17:51 Kiadtisak GEM ADC's Plots
  • 319452 04/25/10 17:34 O.Glamazdin Moller magnet has been filled
  • 319447 04/25/10 17:29 LeRose Hall A Tools
  • 319443 04/25/10 16:40 A. Camsonne Gem HV off
  • 319442 04/25/10 16:33 mcnulty Moved to empty target. VDC HVs off and quartz HV set for integrating mode
  • 319441 04/25/10 16:33 R. Michaels empty target
  • 319438 04/25/10 16:18 R. Michaels more thin Pb runs
  • 319434 04/25/10 16:06 R. Michaels thin lead target running
  • 319431 04/25/10 16:03 A. Camsonne GEM HV increased
  • 319430 04/25/10 16:02 mcnulty online Q2 for Pb#3 at 0.2kHz T5 rate
  • 319426 04/25/10 15:55 Mindy 04/25/2010 Day Shift Summary
  • 319425 04/25/10 15:52 R. Michaels Qsq runs on lead/diamond. Rate scan.
  • 319424 04/25/10 15:51 mcnulty online Q2 for Pb#3 at 3kHz T5 rate
  • 319417 04/25/10 15:41 mcnulty online Q2 for Pb#3 at 9.3kHz T5 rate
  • 319414 04/25/10 15:37 mcnulty online Q2 for Pb#3 at 57kHz T5 rate
  • 319409 04/25/10 15:31 mcnulty online Q2 for Pb#3 at 125kHz T5 rate
  • 319408 04/25/10 15:29 A. Camsonne GEM HV changed
  • 319403 04/25/10 15:26 mcnulty online Q2 for thin Carbon, raster on
  • 319400 04/25/10 15:19 mcnulty Rarm spectra for thin Carbon with raster
  • 319397 04/25/10 15:16 mcnulty Larm spectra (thin Carbon) with raster on
  • 319394 04/25/10 15:11 mcnulty Larm spectra from thin Carbon (no raster)
  • 319393 04/25/10 15:08 mcnulty Rarm Carbon spectra (no raster)
  • 319391 04/25/10 15:03 R. Michaels spot++
  • 319390 04/25/10 15:01 A. Camsonne GEM current HV settings
  • 319387 04/25/10 14:59 mcnulty online Q2 for (9mg) Carbon Target
  • 319384 04/25/10 14:57 R. Michaels Qsq on thin C12, raster off and on.
  • 319383 04/25/10 14:56 mcnulty L and Rarm Quartz signals
  • 319382 04/25/10 14:49 R. Michaels correction: raster is off
  • 319371 04/25/10 14:29 R. Michaels magnet scan and low-I cavities
  • 319365 04/25/10 13:42 A. Camsonne Gen pedestal file updated
  • 319364 04/25/10 13:39 R. Michaels L-HRS network switch power cycled (late entry)
  • 319363 04/25/10 13:36 mcnulty Rarm VDC HV not ramping up
  • 319362 04/25/10 13:29 mcnulty VDCs on and quartz HV at -2000 for all dets
  • 319360 04/25/10 13:19 Bob and Dust Checks of ADC integrate timing
  • 319356 04/25/10 13:14 ZAFAR AHMED Target is empty target
  • 319342 04/25/10 12:58 mcnulty Moving to BeO target
  • 319341 04/25/10 12:57 M Friend Compton DAQ running at 240Hz
  • 319339 04/25/10 12:53 A. Camsonne Spot++ T8 only 2 uA
  • 319335 04/25/10 12:48 Mindy turn the septum on
  • 319331 04/25/10 12:36 A. Camsonne Upper scintillator left arm turned on
  • 319329 04/25/10 12:17 mercado Available sync signals look fine
  • 319328 04/25/10 11:55 mcnulty Switched quartz dets to counting/sampling mode
  • 319327 04/25/10 11:54 mcnulty changed helicty from 120Hz t0 240Hz
  • 319325 04/25/10 11:42 R. Michaels switching to 240 Hz
  • 319324 04/25/10 11:42 ZAFAR AHMED Amplifiers on cavities - No more than 5uA.
  • 319322 04/25/10 11:14 A. Camsonne Gem 5th chamberHV plugged - reminder GEM HV
  • 319320 04/25/10 10:59 Mindy power-cycle ROC-25-->CODA is back
  • 319319 04/25/10 10:16 R. Michaels reboot beamline slow controls
  • 319318 04/25/10 10:12 Mindy [trouble-shooting guide] temperature monitor works again
  • 319317 04/25/10 09:42 R. Michaels Cleanup of adaqfs.
  • 319316 04/25/10 09:28 ZAFAR AHMED Septum magnet is off
  • 319315 04/25/10 09:25 R. Michaels access, empty target
  • 319314 04/25/10 09:21 R. Michaels Run plan: Qsq and other
  • 319313 04/25/10 09:04 Abdurahim Compton DOCP scan
  • 319312 04/25/10 08:27 Mindy Hall-A general tool
  • 319311 04/25/10 08:12 R. Michaels helicity to Asym box unplugged
  • 319309 04/25/10 07:59 Yoomin Owl Shift Summary
  • 319308 04/25/10 07:54 R. Michaels temp readouts need to be fixed
  • 319305 04/25/10 07:35 Yoomin Target IOC reboot
  • 319304 04/25/10 07:04 Yoomin Compton HV trip
  • 319303 04/25/10 06:08 Yoomin Radiation/Beam stripchart
  • 319302 04/25/10 06:05 Yoomin Reduced beam current 40uA
  • 319301 04/25/10 05:58 Yoomin Prompt not perfect yet
  • 319299 04/25/10 05:54 silwal more plots on run 3119, helicity toggled @ 120 Hz
  • 319297 04/25/10 05:27 Yoomin Right Q1 fluctuating
  • 319296 04/25/10 05:14 Yoomin Radiation/Beam stripchart
  • 319295 04/25/10 05:12 Yoomin Prompt works again!
  • 319294 04/25/10 05:11 Yoomin Compton HV trip
  • 319293 04/25/10 05:10 Yoomin Target Screen
  • 319292 04/25/10 05:07 Yoomin Compton HV trip
  • 319289 04/25/10 04:44 Yoomin Compton HV trip/R-Q1 frequent alarm
  • 319288 04/25/10 04:35 Yoomin Hall A Tools Capture
  • 319287 04/25/10 04:28 silwal asym from the 'asym box' measured to be ~ 1.1 %
  • 319286 04/25/10 04:18 Yoomin Compton HV trip
  • 319285 04/25/10 04:07 Yoomin Compton HV trip
  • 319284 04/25/10 03:46 Yoomin Compton HV trip
  • 319281 04/25/10 02:55 Yoomin Compton HV trip
  • 319280 04/25/10 02:49 Yoomin Collimator temperature monitor not working
  • 319272 04/25/10 02:09 Yoomin Temperature monitor not working!
  • 319271 04/25/10 01:55 Yoomin Compton HV trip
  • 319270 04/25/10 01:54 M Friend Compton 120Hz crl fixed
  • 319269 04/25/10 01:39 M Friend /adaql10/work1 full
  • 319266 04/25/10 01:25 Yoomin Rad/Temp/Beam strip chart
  • 319265 04/25/10 01:11 Yoomin Compton HV trip
  • 319264 04/25/10 01:09 Yoomin Target IOC reboot
  • 319263 04/25/10 00:57 Yoomin HV IOC Down.
  • 319261 04/25/10 00:02 bogdanw shift summary
  • Saturday

  • 319258 04/24/10 23:53 silwal 120Hz toggle run
  • 319255 04/24/10 22:11 jhlee prompt *** Break *** segmentation violation
  • 319252 04/24/10 21:32 Abdurahim Compton cavity has been restored
  • 319251 04/24/10 21:26 Mindy comparison of dithering det_vs_bpm slope between 4x4 and 5x5
  • 319248 04/24/10 20:13 jpchen changed alarms
  • 319247 04/24/10 20:07 silwal asym ~1.2 % measured on the scope (asym-box)
  • 319246 04/24/10 19:55 Apar non linearity of LUMI
  • 319245 04/24/10 19:47 jhlee dithering
  • 319244 04/24/10 19:39 bogdanw beam position and radiation monitors
  • 319243 04/24/10 19:34 bogdanw tools and temperatures
  • 319242 04/24/10 19:27 silwal asym measurement from asym box
  • 319241 04/24/10 19:21 bogdanw HV lumi with Pb and 50 muA
  • 319239 04/24/10 19:20 bogdanw HV for Pb at 120 Hz at 50 muA
  • 319237 04/24/10 18:52 ZAFAR AHMED Carbon Target Lumi High Voltages
  • 319235 04/24/10 18:43 bogdanw beam on Pb target
  • 319232 04/24/10 18:40 ZAFAR AHMED Lead Diamond upper Scint. rate at nA.
  • 319224 04/24/10 18:38 ZAFAR AHMED Target is Lead Diamond # 3
  • 319223 04/24/10 18:36 ZAFAR AHMED BeO nA Rate . Slit and Attn. Adjustment.
  • 319222 04/24/10 18:32 ZAFAR AHMED Target is BeO - Beam Recovery
  • 319221 04/24/10 18:26 ZAFAR AHMED Target is Lead Diamond # 3
  • 319220 04/24/10 18:12 ZAFAR AHMED Target is BeO - Beam Recovery
  • 319217 04/24/10 18:03 bogdanw spot++
  • 319216 04/24/10 18:02 adaq spot ++
  • 319209 04/24/10 17:28 ZAFAR AHMED Target is empty target
  • 319208 04/24/10 17:27 A. Camsonne Compton electron detector work
  • 319206 04/24/10 17:20 O.Glamazdin Moller magnet has been filled
  • 319205 04/24/10 17:19 M Friend Compton photon collimator moved
  • 319202 04/24/10 16:29 R. Michaels modified warnings about ADC18
  • 319201 04/24/10 16:24 silwal target: false alarms on gas panels, to do with LD2 that does not exist
  • 319200 04/24/10 16:13 jhlee screenshot of Dithering
  • 319199 04/24/10 16:10 jhlee ROC 23 has the same error meassage as 318478
  • 319197 04/24/10 16:04 Sergey Abrah Day shift summary - 04/24/2010
  • 319194 04/24/10 15:57 R. Michaels Adjusting HV for 120Hz C12
  • 319193 04/24/10 15:34 R. Michaels prep for 120 Hz running
  • 319192 04/24/10 15:31 Apar Changed to 120 Hz
  • 319191 04/24/10 15:28 aserg Helicity frequency changed to 120 Hz
  • 319189 04/24/10 15:13 M Friend Compton running at 120Hz
  • 319188 04/24/10 15:13 R. Michaels info about synch mon status
  • 319186 04/24/10 14:55 aserg No beam
  • 319183 04/24/10 14:42 Bob and Mark check of timing signals in CH
  • 319181 04/24/10 14:23 aserg Beam off
  • 319178 04/24/10 14:05 aserg CODA works again
  • 319176 04/24/10 14:01 aserg CODA not working
  • 319172 04/24/10 13:30 aserg Compton HV trip
  • 319171 04/24/10 13:15 aserg Dithering plot
  • 319170 04/24/10 13:14 aserg Beam and radiation monitors
  • 319169 04/24/10 13:12 aserg Temperature monitor
  • 319166 04/24/10 12:17 adpalmer reboot preformed
  • 319165 04/24/10 12:13 adpalmer target vacuum glitch
  • 319162 04/24/10 11:52 dalton ROOTfiles moved
  • 319161 04/24/10 11:06 aserg Compton HV trip
  • 319151 04/24/10 10:36 aserg Compton HV trip
  • 319148 04/24/10 09:34 aserg Dithering
  • 319145 04/24/10 09:04 aserg Radiation and beam monitors
  • 319144 04/24/10 09:01 aserg Temperature monitor strip
  • 319143 04/24/10 08:47 aserg Compton HV trip
  • 319142 04/24/10 08:33 aserg Compton HV trip
  • 319139 04/24/10 08:05 ZAFAR AHMED April 24 owl sfit summary
  • 319137 04/24/10 07:59 Abdurahim Compton strip chart
  • 319135 04/24/10 07:37 ZAFAR AHMED Flipper worked- Slug no. is 13
  • 319134 04/24/10 07:37 apar Halfwave plate OUT, Slug 13
  • 319133 04/24/10 07:36 ZAFAR AHMED Fipper script error
  • 319131 04/24/10 07:05 Abdurahim Compton HV tripped
  • 319130 04/24/10 07:00 Abdurahim Beam is back. Compton HV tripped.
  • 319129 04/24/10 06:58 Abdurahim beam is off
  • 319126 04/24/10 06:20 Abdurahim Compton HV trip
  • 319125 04/24/10 05:45 ZAFAR AHMED Dithering
  • 319122 04/24/10 05:29 Abdurahim Compton HV trip
  • 319121 04/24/10 05:05 Abdurahim Beam back. Compton HV tripped.
  • 319120 04/24/10 05:01 Abdurahim magnet power supply tripped
  • 319119 04/24/10 05:00 Abdurahim almost two hours with no beam trip
  • 319116 04/24/10 04:50 Abdurahim Target temperature
  • 319115 04/24/10 04:43 Abdurahim Compton rates are beautifully stable
  • 319114 04/24/10 04:41 Abdurahim Compton HV trip
  • 319113 04/24/10 04:32 ZAFAR AHMED Dithering
  • 319110 04/24/10 03:40 ZAFAR AHMED Dithering
  • 319109 04/24/10 03:24 Abdurahim StriptCharts
  • 319097 04/24/10 02:19 Abdurahim Compton HV tripped
  • 319096 04/24/10 01:55 Abdurahim started Compton laser polarization scan
  • 319095 04/24/10 01:32 ZAFAR AHMED Dithering is working
  • 319092 04/24/10 01:28 Abdurahim How to Start Target Strip Chart
  • 319082 04/24/10 00:55 ZAFAR AHMED Dithering is not working
  • 319080 04/24/10 00:28 ZAFAR AHMED Dithering is off
  • 319079 04/24/10 00:27 Abdurahim target epics problem
  • 319078 04/24/10 00:17 ZAFAR AHMED New slug(13) but half wave plate is not changed
  • 319077 04/24/10 00:10 Mindy 04/23/2010 Swing Shift Summary
  • 319076 04/24/10 00:10 R. Michaels putting in thick C12
  • 319075 04/24/10 00:09 N. Muangma target thick carbon
  • Friday

  • 319071 04/23/10 23:49 M Friend Compton photon detector HV
  • 319069 04/23/10 23:43 adaq Compton FADC crl accum thresh changed
  • 319068 04/23/10 23:30 M Friend Compton EPICS reboot
  • 319065 04/23/10 22:29 Mindy temperature become stable now
  • 319062 04/23/10 21:34 Mindy dithering comes back
  • 319058 04/23/10 20:44 Mindy temperature is increasing
  • 319057 04/23/10 20:38 Mindy dithering correlations/slopes (1)
  • 319056 04/23/10 20:23 R.Michaels watch the temps (with bigger raster) !
  • 319055 04/23/10 20:20 Mindy temperature increase?
  • 319054 04/23/10 20:14 M Friend Compton FADC crl accum thresh changed
  • 319053 04/23/10 20:00 M Friend Compton FADC crl change
  • 319048 04/23/10 19:28 M Friend strange Compton EPICS glitch
  • 319047 04/23/10 19:25 Mindy Hall-A general tool
  • 319046 04/23/10 19:06 Mindy dithering works well
  • 319045 04/23/10 19:05 Mindy spot++ for production
  • 319040 04/23/10 18:46 R. Michaels reducing spot size, possible burn ring
  • 319028 04/23/10 18:32 dalton High statistics and zoom of recent Pb/C runs
  • 319024 04/23/10 18:24 R. Michaels Moving beam horizontally
  • 319011 04/23/10 17:49 R. Michaels vertical scan, fishing for Cu frame.
  • 319006 04/23/10 17:36 Mindy FYI: bad dithering run-list
  • 319001 04/23/10 17:32 R. Michaels Pb/D raster plots, bigger raster
  • 319000 04/23/10 17:23 dalton Higher statistics and zoom of previous Pb/D
  • 318997 04/23/10 17:18 R. Michaels Pb/D raster plots, and plan
  • 318992 04/23/10 17:04 M Friend Compton FADC crl change
  • 318991 04/23/10 17:04 N Muangma target PbD#3
  • 318988 04/23/10 17:00 R. Michaels BeO sampling with 4x3 raster
  • 318983 04/23/10 16:45 R. Michaels Raster scan. Spot_L
  • 318975 04/23/10 16:17 mercado Thick Carbon Detector Widths and Correlations
  • 318974 04/23/10 16:16 jwexler HVs changed back.
  • 318973 04/23/10 16:15 N. Muangma target moves BeO
  • 318972 04/23/10 16:04 Mindy Hall-A general tool
  • 318969 04/23/10 16:01 jwexler Shift summary - Day 04/23/2010
  • 318967 04/23/10 15:56 jwexler Change to thick carbon HVs.
  • 318965 04/23/10 15:48 M Friend Compton FADC crl accum thresh changed
  • 318964 04/23/10 15:41 jwexler Changed target - thick carbon
  • 318963 04/23/10 15:39 jwexler Detector and Lumi HVs changed
  • 318961 04/23/10 15:30 O.Glamazdin Moller magnet has been filled
  • 318958 04/23/10 14:18 aserg Dithering
  • 318955 04/23/10 13:37 jwexler LeCroy beamline crate alarm
  • 318952 04/23/10 12:56 Abdurahim Compton cavity and laser is runnning again
  • 318951 04/23/10 12:09 jwexler Hall A tools
  • 318950 04/23/10 12:08 jwexler Target screen and stripcharts
  • 318947 04/23/10 12:03 mercado Re-prompting of slug 10 in progress
  • 318946 04/23/10 11:49 mercado Re-Prompting of slug 11 in progress
  • 318945 04/23/10 11:48 Abdurahim mysterious compton epics reboot
  • 318944 04/23/10 11:36 mercado Wrong BCM mapping.
  • 318943 04/23/10 11:36 dalton Changes to database
  • 318940 04/23/10 11:01 aserg Dithering
  • 318939 04/23/10 11:01 jwexler Radiation and temp. stripcharts
  • 318938 04/23/10 10:42 mercado Changed Gains of BPM 18-bit ADCs
  • 318924 04/23/10 08:36 jwexler Compton HV trip
  • 318923 04/23/10 08:25 Yoomin Compton HV trip
  • 318920 04/23/10 07:52 Yoomin PREX DAQ crash
  • 318915 04/23/10 07:20 Yoomin Target stripchart
  • 318914 04/23/10 07:18 Yoomin Compton HV trip
  • 318911 04/23/10 06:39 Yoomin Radiation/Temperature/Beam
  • 318910 04/23/10 06:38 Yoomin Hall A Tools
  • 318909 04/23/10 06:35 Yoomin Compton HV trip
  • 318908 04/23/10 06:23 ZAFAR AHMED Dithering
  • 318907 04/23/10 06:17 Yoomin Compton Run 22654 ended due to HWP flip
  • 318905 04/23/10 06:14 ZAFARA AHMED New Slug - Half Wave Plate IN
  • 318904 04/23/10 06:14 apar Halfwave plate IN, Slug 12
  • 318900 04/23/10 05:42 Yoomin Target IOC reboot
  • 318898 04/23/10 05:13 ZAFAR AHMED Dithering
  • 318897 04/23/10 05:10 Yoomin Target Screen Capture
  • 318895 04/23/10 05:09 Yoomin Compton Strip Chart
  • 318892 04/23/10 04:10 ZAFAR AHMED Dithering
  • 318891 04/23/10 03:59 Yoomin HV trip
  • 318890 04/23/10 03:36 ZAFARA AHMED Ditheirng is working
  • 318889 04/23/10 03:35 ZAFARA AHMED Dithering error(2)
  • 318888 04/23/10 03:33 ZAFARA AHMED Dithering error
  • 318885 04/23/10 02:40 Yoomin Compton HV trip
  • 318875 04/23/10 02:13 M Friend Compton EPICS rate changed
  • 318874 04/23/10 02:08 M Friend Compton prescale set to 32
  • 318873 04/23/10 01:51 ZAFAR AHMED Dithering
  • 318872 04/23/10 01:44 M Friend Compton photon beam position
  • 318870 04/23/10 01:38 Yoomin Target screen capture - beam recovered
  • 318861 04/23/10 00:44 A. Camsonne Electron detector analog signal - scope setting
  • 318860 04/23/10 00:29 R. Michaels upper scint HV off
  • 318859 04/23/10 00:29 A. Camsonne Electron detector analog signal
  • 318857 04/23/10 00:28 R. Michaels raster check on Pb/D
  • 318853 04/23/10 00:19 Yoomin Target Pb Diamond #3
  • 318852 04/23/10 00:19 R. Michaels BeO check, establish few nA
  • 318851 04/23/10 00:09 Yoomin Target motion to BeO
  • 318849 04/23/10 00:06 R. Michaels spot++ and beam position
  • Thursday

  • 318844 04/22/10 23:46 jpchen swing shift summary
  • 318843 04/22/10 23:37 R. Michaels turned on septum
  • 318842 04/22/10 23:32 Moller test run finished
  • 318841 04/22/10 23:30 O.Glamazdin Moller measurements result
  • 318839 04/22/10 23:15 M Friend Compton DAQ set to 240Hz
  • 318838 04/22/10 23:11 O.Glamazdin Helicity board settins have been restored
  • 318833 04/22/10 22:47 A. Camsonne GEM pedestal runs
  • 318824 04/22/10 22:16 M Friend Compton FADC crl far accum thresh changed
  • 318811 04/22/10 18:07 dalton Energy excursion last night
  • 318805 04/22/10 16:53 jpchen target ioc rebooted
  • 318803 04/22/10 16:42 jpchen target temperature reading stopped since this morning
  • 318802 04/22/10 16:06 J. Zhang Shift Summary
  • 318800 04/22/10 15:38 M Friend Compton DAQ set to 30Hz
  • 318799 04/22/10 15:37 M Friend Compton FADC crl change
  • 318798 04/22/10 15:06 jwexler Tiefenback dp/p
  • 318797 04/22/10 15:02 O.Glamazdin Helicity Board changes
  • 318796 04/22/10 14:30 M Friend Compton FADC crl accum thresh changed
  • 318793 04/22/10 14:15 jwexler Target screen/stripchart
  • 318792 04/22/10 14:12 jwexler Temp/Rad charts
  • 318791 04/22/10 14:09 jwexler Hall A Tools
  • 318790 04/22/10 13:59 M Friend Compton DAQ restored
  • 318789 04/22/10 13:48 J. Zhang Moller VME crates rebooted
  • 318788 04/22/10 13:39 jwexler Back to beam permit
  • 318787 04/22/10 13:39 A. Camsonne Did not receive page in the HRS hut
  • 318785 04/22/10 13:29 A. Camsonne GEM gas line and electronics hooked up on fifth chamber
  • 318784 04/22/10 13:19 J. Zhang Moller solenoid was refilled
  • 318783 04/22/10 13:18 Moller solenoid was refilled
  • 318775 04/22/10 12:42 M Friend Compton photon signal in the CH
  • 318760 04/22/10 10:33 jwexler Septum off
  • 318759 04/22/10 10:23 ZAFAR AHMED Going to controlled access
  • 318756 04/22/10 09:59 M Friend Compton DAQ set to take pulser runs
  • 318755 04/22/10 09:56 M Friend Compton table moved
  • 318753 04/22/10 09:35 M Friend Cavity flip script
  • 318752 04/22/10 09:33 M Friend Compton cavity locking
  • 318747 04/22/10 09:18 jwexler Target moved to empty & Error in GUI handler
  • 318745 04/22/10 09:08 ZAFAR AHMED Beam study starts.
  • 318743 04/22/10 09:04 M Friend Compton epics reboot
  • 318742 04/22/10 08:49 aserg Wrong HWP info
  • 318739 04/22/10 08:37 jwexler Compton cavity not locked
  • 318738 04/22/10 08:15 aserg Dithering
  • 318737 04/22/10 08:05 Yoomin Owl Shift Summary
  • 318733 04/22/10 07:08 Yoomin Compton Background Rate
  • 318730 04/22/10 06:46 Yoomin HV trip
  • 318729 04/22/10 06:31 Yoomin Target strip chart
  • 318728 04/22/10 06:26 Yoomin Stripchart radiation/temperature/beam
  • 318727 04/22/10 06:20 Yoomin HV trip
  • 318724 04/22/10 06:10 Yoomin Compton crash again
  • 318723 04/22/10 05:30 Yoomin Compton strip chart
  • 318722 04/22/10 05:26 Hisham Coda Crash
  • 318714 04/22/10 03:01 Yoomin Compton recovered
  • 318713 04/22/10 02:53 Yoomin Compton DAQ crash
  • 318704 04/22/10 02:14 Yoomin From Compton run 22604, IHWP is OUT
  • 318703 04/22/10 02:09 Hisham CH scaler GUI
  • 318702 04/22/10 02:01 Yoomin Target Strip Chart
  • 318701 04/22/10 02:00 Yoomin Target Screen capture
  • 318700 04/22/10 01:58 Yoomin Please refer to ELOG 1520004
  • 318698 04/22/10 01:53 Yoomin dpp
  • 318696 04/22/10 01:43 Yoomin Radiation/Temperature/BPMs
  • 318695 04/22/10 01:38 Yoomin Hall A Tools
  • 318694 04/22/10 01:33 Yoomin HV trip
  • 318692 04/22/10 01:00 Hisham HAPPEX
  • 318691 04/22/10 00:59 apar Halfwave plate OUT, Slug 11
  • 318686 04/22/10 00:05 V. Sulkosky Swing Shift Summary
  • Wednesday

  • 318685 04/21/10 23:27 V. Sulkosky dpp Excursion
  • 318682 04/21/10 23:12 adaq shift crews
  • 318681 04/21/10 23:08 N Muangma empty plots
  • 318679 04/21/10 22:25 Abdurahim Compton cavity lock has been restored
  • 318677 04/21/10 22:18 gjin hall a tools screen
  • 318676 04/21/10 22:17 gjin target screen
  • 318675 04/21/10 22:00 V. Sulkosky Compton HV Changed
  • 318674 04/21/10 21:56 M Friend Glitch in FADC Compton data
  • 318671 04/21/10 21:22 M Friend Compton EPICS trigger rate change
  • 318670 04/21/10 21:19 V. Sulkosky Compton HV Change
  • 318669 04/21/10 21:17 V. Sulkosky Collimator Temperatures
  • 318668 04/21/10 21:11 N Muangma BCM signal jump
  • 318665 04/21/10 20:26 adaq Compton HV Trips (1st Half)
  • 318662 04/21/10 19:51 N Muangma dither plots
  • 318661 04/21/10 19:19 V. Sulkosky Radiation Strip Chart
  • 318658 04/21/10 18:57 N Muangma double peak structure
  • 318651 04/21/10 18:31 V. Sulkosky Compton Stripchart
  • 318648 04/21/10 18:10 silwal slug 9 transmission plots
  • 318647 04/21/10 17:58 silwal slug 6 transmission plots
  • 318646 04/21/10 17:49 V. Sulkosky Hall A Tools
  • 318645 04/21/10 17:46 silwal slug 7 transmission plots
  • 318644 04/21/10 17:37 M Friend Compton discriminator set high
  • 318641 04/21/10 17:05 M Friend Compton rates with slits closed
  • 318637 04/21/10 16:02 M Friend Compton EPICS variable
  • 318636 04/21/10 15:58 J. Zhang HallA Tools Screen
  • 318634 04/21/10 15:25 cusanno Target Screenshots
  • 318632 04/21/10 14:52 silwal slug 8 transmission plots
  • 318627 04/21/10 13:05 cusanno Target Pressure when beam is back
  • 318625 04/21/10 12:33 O.Glamazdin Moller magnet has been filled
  • 318624 04/21/10 12:28 cusanno To do before rebooting target IOC
  • 318623 04/21/10 11:53 J. Zhang Control access
  • 318622 04/21/10 11:51 J. Zhang Beam was downdue to the injector problem
  • 318621 04/21/10 11:47 adaq Beam was not stable for run 3009
  • 318619 04/21/10 11:01 cusanno Target Alarm Handler
  • 318618 04/21/10 10:47 cusanno Target Alarms as in entry 318236
  • 318617 04/21/10 10:43 O.Glamazdin Slug 10 for IHWP=IN
  • 318615 04/21/10 10:38 apar Halfwave plate IN, Slug 10
  • 318614 04/21/10 10:38 Jixie Zhang Half Plate changed
  • 318613 04/21/10 10:36 silwal Dithering response changing
  • 318606 04/21/10 10:33 cusanno Target IOC
  • 318604 04/21/10 10:25 Abdurahim Compton laser polarization scan (RIGHT) is completed
  • 318603 04/21/10 10:15 O.Glamazdin Radiation level in the Hall
  • 318600 04/21/10 09:55 O.Glamazdin Dithering looks strange
  • 318599 04/21/10 09:23 J. Zhang Day Shift Summary (8:00-16:00)
  • 318596 04/21/10 08:43 jixie Compton HV(card11, channel 1) tripped
  • 318594 04/21/10 08:23 Abdurahim Owl shift summary
  • 318591 04/21/10 07:54 Abdurahim radiation level OK ?
  • 318590 04/21/10 07:53 Abdurahim Now dithering can be seen in BPM strip chart
  • 318588 04/21/10 07:29 Abdurahim Compton background is getting worse
  • 318586 04/21/10 07:25 Abdurahim Target gui
  • 318585 04/21/10 07:24 Abdurahim Target temperature
  • 318584 04/21/10 07:16 Abdurahim Hall A general tools
  • 318583 04/21/10 07:15 Abdurahim BPM & BCM
  • 318582 04/21/10 07:14 Abdurahim radiation monitor
  • 318581 04/21/10 07:13 Abdurahim collimator temperatures
  • 318580 04/21/10 06:52 Abdurahim Compton PMT HV stripped
  • 318579 04/21/10 06:45 Abdurahim Dithering cannot be seen in the BPM strip chart
  • 318578 04/21/10 06:44 Hisham Dithering
  • 318573 04/21/10 05:37 Abdurahim Compton PMT HV trip
  • 318572 04/21/10 05:23 Hisham Blumi Signals Around
  • 318571 04/21/10 05:06 Abdurahim Target screen
  • 318570 04/21/10 05:05 Abdurahim Target temperature and pressure
  • 318567 04/21/10 04:30 Abdurahim Compton strip chart during laser polarization scan
  • 318564 04/21/10 03:50 Abdurahim Radiation level a concern
  • 318553 04/21/10 01:44 Abdurahim Hall A general tools
  • 318552 04/21/10 01:42 Abdurahim Target gui
  • 318551 04/21/10 01:41 Abdurahim Target temperature
  • 318550 04/21/10 01:40 Abdurahim radiation monitor
  • 318549 04/21/10 01:39 Abdurahim BCM and BPM strip chart
  • 318548 04/21/10 01:38 Abdurahim Collimator temperatures
  • 318545 04/21/10 01:01 M Friend Compton HV to -1700
  • 318542 04/21/10 00:12 V. Sulkosky Swing Shift Summary (better format)
  • 318541 04/21/10 00:03 V. Sulkosky Swing Shift Summary
  • Tuesday

  • 318539 04/20/10 23:57 V. Sulkosky Collimator Temperatures
  • 318534 04/20/10 23:34 apar Halfwave plate OUT, Slug 9
  • 318531 04/20/10 23:28 V. Sulkosky Re: Recalibrated BCMs
  • 318527 04/20/10 23:13 V. Sulkosky Left Q1 current tripped (TECHs)
  • 318526 04/20/10 22:54 V. Sulkosky Compton Stripchart
  • 318525 04/20/10 22:54 mcanan Target Screens
  • 318524 04/20/10 22:41 V. Sulkosky Compton GSO HV to -1.65 kV
  • 318521 04/20/10 22:23 R. Michaels re: opmanual
  • 318520 04/20/10 22:08 M Friend Compton op-manual
  • 318519 04/20/10 21:54 M Friend Compton rates with target in
  • 318518 04/20/10 21:48 V. Sulkosky Radiation Monitor Stripchart
  • 318517 04/20/10 21:48 jhlee screenshot for PANGUIN
  • 318513 04/20/10 21:24 V. Sulkosky Raster Check on Lead/Diamond
  • 318511 04/20/10 21:21 V. Sulkosky Rate on Lead/Diamond
  • 318509 04/20/10 21:16 mcanan Target Motion
  • 318508 04/20/10 21:15 V. Sulkosky Rate check on BeO
  • 318507 04/20/10 21:09 M Friend Compton rates after tune
  • 318506 04/20/10 21:08 mcanan Target Motion
  • 318505 04/20/10 21:01 V. Sulkosky Spot++ Check 17 x 7 Empty
  • 318498 04/20/10 20:10 V. Sulkosky Beam Restored to Hall A
  • 318495 04/20/10 19:19 jhlee Run 2985 has "end fail"
  • 318491 04/20/10 19:00 V. Sulkosky Hall A Tools
  • 318490 04/20/10 18:51 V. Sulkosky Left Q1 off and Regulators off
  • 318489 04/20/10 18:40 V. Sulkosky BCM Logger Updated
  • 318482 04/20/10 18:29 jhlee Roc 23 restarted
  • 318481 04/20/10 18:29 R. Michaels explanation of ROC23 errors
  • 318478 04/20/10 17:49 jhlee the right error message at ROC 23 with run 2983
  • 318477 04/20/10 17:27 V. Sulkosky Beam Ready but ...
  • 318476 04/20/10 17:19 jwexler ROC23 power cycled.
  • 318472 04/20/10 16:35 V. Sulkosky South LINAC Vacuum Valve
  • 318470 04/20/10 16:34 R. Michaels swing shift activities
  • 318467 04/20/10 16:03 Ole Hansen Shift Summary
  • 318466 04/20/10 15:47 Fansler/Sega RQ1/RQ2
  • 318463 04/20/10 15:19 Ole Hansen Cycling R-HRS Q2/Q3
  • 318462 04/20/10 14:59 Ole Hansen Septum ON
  • 318461 04/20/10 14:37 A. Camsonne GEM activities
  • 318460 04/20/10 14:19 Ole Hansen Access activities
  • 318459 04/20/10 14:09 O.Glamazdin Moller magnet has been filled
  • 318458 04/20/10 13:59 A. Camsonne GEM INFN HV
  • 318457 04/20/10 13:58 Abdurahim a new webcam for Compton quarter wave plate
  • 318456 04/20/10 13:56 test test
  • 318450 04/20/10 10:38 M Friend Compton epics reboot
  • 318443 04/20/10 10:28 M Friend Compton PMT base current
  • 318441 04/20/10 10:22 R. Michaels Compton DAQ tests, notes
  • 318439 04/20/10 09:55 R. Michaels Compton DAQ configs
  • 318438 04/20/10 09:43 paschke runs require re-prompting
  • 318435 04/20/10 09:08 Jixie Zhang HV tripped: card11_channel1 (compton)
  • 318429 04/20/10 07:56 ZAFAR AHMED April 20 Owl Shift Summary
  • 318427 04/20/10 07:45 cusanno Empty Target IN
  • 318425 04/20/10 07:40 cusanno Target Screenshot
  • 318420 04/20/10 06:22 ZAFAR AHMED Production
  • 318419 04/20/10 05:47 cusanno Target Screenshots (contin'd)
  • 318418 04/20/10 05:44 cusanno Target Screenshot
  • 318417 04/20/10 05:43 efuchey Dithering ON
  • 318415 04/20/10 05:38 ZAFAR AHMED Beam recovery 5uA. Raster
  • 318414 04/20/10 05:35 efuchey other errata
  • 318406 04/20/10 05:13 ZAFAR AHMED Beam Recover. Not right raster at lead target 5nA.
  • 318398 04/20/10 05:02 ZAFAR AHMED Beam Recovery-5nA at target Pb Diamond 3
  • 318397 04/20/10 04:58 cusanno Lead Target IN
  • 318394 04/20/10 04:55 ZAFAR AHMED Beam Recovery-5 nA-BeO
  • 318392 04/20/10 04:41 cusanno BeO Target IN
  • 318391 04/20/10 04:38 ZAFAR AHMED spot++ Beam recovry 17x7 Raster
  • 318388 04/20/10 04:33 ZAFAR AHMED Spot++ Beam recovery
  • 318387 04/20/10 04:33 efuchey errata
  • 318384 04/20/10 04:16 ZAFAR AHMED Beam is back.
  • 318383 04/20/10 04:13 cusanno Target Screenshot
  • 318380 04/20/10 03:36 ZAFAR AHMED Beam update
  • 318377 04/20/10 02:53 ZAFAR AHMED No beam - Optics on call is called by MCC
  • 318376 04/20/10 02:52 cusanno Empty Target
  • 318375 04/20/10 02:37 M Friend Running the Compton DAQ tonight
  • 318366 04/20/10 02:04 ZAFAR AHMED No beam for last one hour.
  • 318363 04/20/10 01:22 ZAFAR AHMED No beam for last 20 minutes.
  • 318360 04/20/10 01:12 ZAFAR AHMED No beam for last 10 minutes.
  • 318355 04/20/10 01:00 ZAFAR AHMED Spot++
  • 318352 04/20/10 00:48 ZAFAR AHMED Dithring coils at the beginning of owl shift.
  • 318351 04/20/10 00:46 ZAFAR AJMED Parity and fast feedback at the beginning fo owl shift
  • 318348 04/20/10 00:34 ZAFAR AHMED Compton rates chart - start of owl shift
  • 318347 04/20/10 00:29 ZAFAR AHMED Upstream collimator Current charts Owl shift starts
  • 318346 04/20/10 00:25 ZAFAR AHMED Hall A Tools at the beginning of owl shift
  • 318342 04/20/10 00:07 Megan F. / M 04/19/2010 Swing shift summary
  • Monday

  • 318341 04/19/10 23:40 Yoomin Target Stat during the beam recovery procedure
  • 318339 04/19/10 23:29 Mindy Hall-A general tool (beam recovery procedure)
  • 318338 04/19/10 23:21 Yoomin Target Pb Diamond #3
  • 318337 04/19/10 23:17 Mindy spot++ check (beam recovery procedure)
  • 318332 04/19/10 22:52 Bob/Kent reset RHRS Q2, leave local mode (TECHS)
  • 318331 04/19/10 22:52 Yoomin Target motion to Empty
  • 318330 04/19/10 22:43 Mindy right-arm Q2 recycling
  • 318327 04/19/10 22:29 O.Glamazdin Moller magnet has been filled
  • 318326 04/19/10 22:09 Yoomin JT Valve
  • 318325 04/19/10 21:54 Yoomin Target chart after IOC reboot
  • 318324 04/19/10 21:52 Yoomin Target IOC rebooted
  • 318322 04/19/10 21:31 Yoomin Target IOC Frozen?
  • 318319 04/19/10 20:20 Yoomin Target Screen Capture - CryoStat
  • 318318 04/19/10 20:19 Yoomin Target Screen Capture
  • 318317 04/19/10 20:15 R. Michaels suggest NO change to JT valve
  • 318315 04/19/10 19:57 O.Glamazdin PREX CODA crush
  • 318312 04/19/10 19:41 Ole Hansen New 2TB disks
  • 318311 04/19/10 19:36 paschke DAQ hang; ROC31 doesn't boot, moved to PrexNoQweak configuration
  • 318299 04/19/10 17:40 M Friend Momentary left magnet alarm
  • 318298 04/19/10 17:37 Yoomin JT2 closed one step more
  • 318297 04/19/10 17:14 dalton batt2 corrected in past datamaps to run 2800
  • 318296 04/19/10 17:14 paschke Feedback PITA slope
  • 318295 04/19/10 17:04 dalton lobeam changed to 90000 channels on bcm3
  • 318294 04/19/10 16:59 paschke correction: pita slope from injector studies at different RHWP position
  • 318292 04/19/10 16:58 mercado Added batt variables to db files for crosstalk tests
  • 318290 04/19/10 16:49 silwal dithering response much worse after quad changes for compton
  • 318289 04/19/10 16:26 Yoomin JT valve one step closed
  • 318288 04/19/10 16:17 M Friend Dithering is ON
  • 318286 04/19/10 16:07 Abdurahim Target temperature
  • 318281 04/19/10 14:59 mcnulty Quartz det pulse height dists before and after s0 removal
  • 318278 04/19/10 13:51 M Friend Compton photon detector can see the target
  • 318277 04/19/10 13:44 aserg spot++ after compton
  • 318272 04/19/10 13:20 Abdurahim Target -> Pb Diamond#3
  • 318271 04/19/10 13:10 Abdurahim target -> Empty
  • 318268 04/19/10 13:00 aserg changing in flipper
  • 318267 04/19/10 12:49 Abdurahim Target -> Pb Diamond #3
  • 318266 04/19/10 12:46 M Friend Compton interaction point
  • 318265 04/19/10 12:44 M Friend Compton rates after tune
  • 318264 04/19/10 12:16 Ole Hansen Computer work planned
  • 318263 04/19/10 12:14 R. Michaels adaql10 problems
  • 318262 04/19/10 12:11 mercado adaql10 down
  • 318259 04/19/10 11:37 M Friend Compton tune started
  • 318258 04/19/10 11:24 aserg DITHERING off
  • 318257 04/19/10 11:11 Abdurahim Target -> Empty
  • 318253 04/19/10 10:44 mercado LUMI 1 HV was off.
  • 318246 04/19/10 10:19 Abdurahim Target strip chart
  • 318244 04/19/10 10:13 R. Michaels reboot iocha11 and 22
  • 318243 04/19/10 10:13 Abdurahim Target motion
  • 318242 04/19/10 10:04 mercado Added battery to datamap
  • 318236 04/19/10 08:48 Abdurahim target
  • 318232 04/19/10 08:00 ZAFAR AHMED 04/19/10 Owl Shift Summary
  • 318231 04/19/10 07:40 cusanno Target Temperature and Alarm Handler
  • 318230 04/19/10 07:31 cusanno Target IOC rebooted
  • 318229 04/19/10 07:25 cusanno Target Alarm Handler frozen?
  • 318226 04/19/10 07:15 cusanno LeCroy HV Crate rebooted
  • 318225 04/19/10 07:04 ZAFAR AHMED Hall A tools
  • 318224 04/19/10 06:59 ZAFAR AHMED Lecroy HV is frozen
  • 318223 04/19/10 06:42 ZAFAR AHMED spot++
  • 318218 04/19/10 06:08 cusanno Target Pressure Readout
  • 318215 04/19/10 06:00 cusanno Target Screenshot
  • 318210 04/19/10 03:44 ZAFAR AHMED Record Runs of PREx
  • 318206 04/19/10 03:11 cusanno Target Screenshots
  • 318197 04/19/10 02:25 ZAFAR AHMED No beam drop for last two hours
  • 318196 04/19/10 01:53 ZAFAR AHMED No beam drop for last one hour
  • 318193 04/19/10 01:24 efuchey Lack of disk space for data replay again -> Solved by logging in and out
  • 318192 04/19/10 01:06 silwal PAN multiplet analyzer sick or bmw variables not setup correctly??
  • 318191 04/19/10 01:03 efuchey Lack of disk space for data replay again
  • 318190 04/19/10 00:56 ZAFAR AHMED HACB_HPE1313A = 4.908
  • 318189 04/19/10 00:55 cusanno target temperature 21.5 K
  • 318188 04/19/10 00:53 ZAFAR AHMED HACB_HPE1313A = 4.88818
  • 318187 04/19/10 00:51 ZAFAR AHMED HAC_HPE1313A is 4.87842
  • 318184 04/19/10 00:47 dalton lobeam extent lengthened for runs 2901-2928
  • 318183 04/19/10 00:46 ZAFAR AHMED HACB_HPE1313A close to 5 Watts
  • 318181 04/19/10 00:12 apar Halfwave plate OUT, Slug 7
  • 318180 04/19/10 00:04 M Friend Compton background rates
  • 318179 04/19/10 00:02 LeRose Shift Summary
  • 318178 04/19/10 00:00 silwal added 'Scaler_bad' cuts to .db files for run 2815-2926
  • Sunday

  • 318176 04/18/10 23:56 LeRose IHWP OUT
  • 318175 04/18/10 23:56 Yoomin Target Chart
  • 318173 04/18/10 23:21 LeRose Production running resumed
  • 318171 04/18/10 23:05 mcnulty raster current scan
  • 318168 04/18/10 22:58 Yoomin Target to Pb Diamond #3
  • 318166 04/18/10 22:53 Yoomin Target BeO
  • 318165 04/18/10 22:48 mcnulty raster size check
  • 318162 04/18/10 22:35 Yoomin Empty target
  • 318161 04/18/10 22:20 dalton Current ramp-up rate less than half after Moller
  • 318160 04/18/10 22:17 silwal Added 'Scaler_bad' cut to the .db files
  • 318159 04/18/10 22:02 mcnulty VDCs, upper scint turned off
  • 318158 04/18/10 21:54 mcnulty called for access to switch to integrating mode
  • 318157 04/18/10 21:54 jwexler Right A_T detector moved.
  • 318154 04/18/10 21:45 mcnulty Rarm Quartz y alignment check after s0 removal
  • 318153 04/18/10 21:44 mcnulty Rarm Quartz x alignment after s0 removal
  • 318152 04/18/10 21:42 mcnulty Quartz y alignment check after s0 removal
  • 318151 04/18/10 21:41 mcnulty Quartz x alignment check after s0 removal
  • 318138 04/18/10 21:19 Yoomin Target to THIN Carbon
  • 318137 04/18/10 21:17 mcnulty spot check
  • 318135 04/18/10 21:00 mcnulty Access activity (s0's out)
  • 318134 04/18/10 20:56 Yoomin Target to Empty
  • 318133 04/18/10 20:48 LeRose Septum is On
  • 318132 04/18/10 19:05 R. Michaels tentative conclusions: Qsq, S0, etc
  • 318131 04/18/10 19:03 mcnulty Q2 for s0 trigger at lower current
  • 318130 04/18/10 18:57 R. Michaels ask for access
  • 318129 04/18/10 18:51 benesch optics results of Compton background test 1438-1455
  • 318128 04/18/10 18:49 R. Michaels Qsq measurements, better now
  • 318125 04/18/10 18:46 mcnulty Q2 for s0 trigger
  • 318104 04/18/10 18:28 mcnulty Q2 for Upper Scint trigger
  • 318103 04/18/10 18:25 Yoomin Target at Pb Diamond #3
  • 318100 04/18/10 18:23 R. Michaels correction: Pb/diamond
  • 318099 04/18/10 18:23 R. Michaels correction: Pb/diamond target #3
  • 318098 04/18/10 18:16 A. Camsonne GEM HV off
  • 318095 04/18/10 18:15 Abdurahim Compton Fiber Amp turned back on
  • 318093 04/18/10 18:13 R. Michaels Podd database problems
  • 318092 04/18/10 18:12 mcnulty L and Rarm Q2 plots, upper Scint trigger (quick and dirty)
  • 318086 04/18/10 18:08 mcnulty Larm alignment plots for thick Lead
  • 318085 04/18/10 18:06 mcnulty Rarm alignment on Thick Pb
  • 318082 04/18/10 17:56 mcnulty quartz counting mode ADC spectra
  • 318081 04/18/10 17:53 R. Michaels rates with super thin C12
  • 318080 04/18/10 17:52 A. Camsonne GEM HV lowered to 4100 V
  • 318077 04/18/10 17:49 A. Camsonne GEM INFN top chamber
  • 318064 04/18/10 17:36 R. Michaels had septum off (oops !)
  • 318063 04/18/10 17:34 M Friend Compton epics reboot
  • 318062 04/18/10 17:31 Yoomin Target at THIN Carbon
  • 318060 04/18/10 17:26 R. Michaels Qsq measurements, part 1
  • 318046 04/18/10 16:49 R. Michaels strange background with beam off
  • 318045 04/18/10 16:45 Yoomin Target moved to Pb Diamond #3
  • 318044 04/18/10 16:40 R. Michaels establishing few-nA on BeO
  • 318043 04/18/10 16:26 R. Michaels visual BeO check
  • 318042 04/18/10 16:15 Yoomin Target moved to BeO
  • 318039 04/18/10 16:09 R. Michaels spot++ checks.
  • 318037 04/18/10 16:04 M Friend Compton rates during quad changes
  • 318032 04/18/10 15:44 R. Michaels HRS DAQs revived
  • 318029 04/18/10 15:24 A. Camsonne Beamline HV crate power cycled
  • 318028 04/18/10 15:21 ZAFAR AHMED Integrating to counting mode
  • 318027 04/18/10 15:04 A. Camsonne GEM HV on
  • 318026 04/18/10 14:37 Mindy turn the septum off
  • 318025 04/18/10 14:33 hisham Dithering off
  • 318024 04/18/10 14:32 Mindy move the target to the empty
  • 318022 04/18/10 14:16 Jack Gas leak in BigBite ethanol bubbler
  • 318017 04/18/10 13:13 Mindy gas leak warning
  • 318016 04/18/10 12:47 mercado Detector Voltages Updated on Opsmanual
  • 318015 04/18/10 12:31 mercado New Detector Voltages for PREX detectors
  • 318010 04/18/10 12:20 mercado Detector Voltages Before Tweak
  • 318009 04/18/10 12:18 Mindy nobody knows why right-arm HV was turned off
  • 318007 04/18/10 12:15 mercado LUMI Pedestals - Run 2913
  • 318005 04/18/10 12:06 hisham Dithering
  • 318004 04/18/10 12:02 Mindy reboot ROC23 in the CH to turn dithering on
  • 318003 04/18/10 12:01 mercado Looking at LUMI Saturation
  • 318002 04/18/10 11:59 hisham Right arm detector output
  • 318001 04/18/10 11:59 Guy Ron Target Vacuum not Logged
  • 318000 04/18/10 11:55 M. Jen reboot right-arm HV setting
  • 317989 04/18/10 10:28 D. Parno LHRS DAQ Problems
  • 317986 04/18/10 09:46 paschke IHWP flip procedure, Run 2908 charge asymmetry
  • 317983 04/18/10 08:30 R. Michaels disk cleanup
  • 317982 04/18/10 08:29 Mindy Target
  • 317981 04/18/10 08:13 efuchey Target temperature
  • 317978 04/18/10 08:03 cusanno Owl Shift Summary
  • 317976 04/18/10 07:57 efuchey Targets GUI screenshots
  • 317975 04/18/10 07:17 Guy Ron Correction for new location of files
  • 317972 04/18/10 07:11 Guy Ron Dirty hack for disk space problem
  • 317971 04/18/10 06:48 cusanno adaql3 disk full with replayed data
  • 317970 04/18/10 06:31 cusanno Beam back
  • 317968 04/18/10 06:23 cusanno Beam OFF
  • 317966 04/18/10 06:11 cusanno again problem with data replay
  • 317965 04/18/10 05:54 cusanno Data replay back
  • 317964 04/18/10 05:46 cusanno Compton HV OFF
  • 317961 04/18/10 05:23 cusanno Hall A Beamline alarm
  • 317958 04/18/10 04:52 cusanno Problems with data replay
  • 317955 04/18/10 04:14 cusanno Hall A Tools screenshot
  • 317954 04/18/10 04:13 cusanno Temperature and radiation screenshots
  • 317953 04/18/10 04:03 cusanno target screenshot
  • 317951 04/18/10 03:33 cusanno IHWP IN
  • 317941 04/18/10 02:30 cusanno Temperatures and radiation stable
  • 317938 04/18/10 01:31 cusanno Target T/P OK
  • 317937 04/18/10 01:28 cusanno Dithering
  • 317934 04/18/10 00:58 silwal dithering turned on, working fine now
  • 317930 04/18/10 00:06 Guy Ron Temprature Log Window
  • 317929 04/18/10 00:03 Franklin Odd dithering results
  • 317928 04/18/10 00:01 Ole Hansen Swing Shift Summary
  • Saturday

  • 317927 04/17/10 23:57 Guy Ron Recalibrated BCMs
  • 317926 04/17/10 23:56 D. Parno Target tools at 50 uA
  • 317925 04/17/10 23:54 Ole Hansen 50 uA - start production
  • 317922 04/17/10 23:39 Franklin Starting Production Runs at 50 microamps
  • 317921 04/17/10 23:38 Ole Hansen Ramping to 50 uA
  • 317920 04/17/10 23:34 Franklin Dithering turned on
  • 317919 04/17/10 23:32 Ole Hansen Spot seems fine
  • 317918 04/17/10 23:30 Ole Hansen spot++ right for Pb/diamond
  • 317916 04/17/10 23:24 Ole Hansen Right upper scint on
  • 317913 04/17/10 23:20 Ole Hansen Rate on Pb/diamond
  • 317912 04/17/10 23:17 Ole Hansen Pb/diamond target
  • 317911 04/17/10 23:17 D. Parno Moved target to Pb Diamond #3
  • 317910 04/17/10 23:15 Ole Hansen Established "5nA"
  • 317909 04/17/10 22:57 D. Parno Moved target to BeO viewer
  • 317907 04/17/10 22:52 Ole Hansen spot++
  • 317905 04/17/10 22:44 Ole Hansen spot++
  • 317902 04/17/10 22:37 O.Glamazdin Moller results
  • 317901 04/17/10 22:14 Ole Hansen Recovery in progress
  • 317900 04/17/10 21:48 Ole Hansen Moller done, restoring production
  • 317893 04/17/10 18:09 D. Parno Target tools snapshot
  • 317892 04/17/10 17:58 O.Glamazdin BPM IPM1H01 reading is noisy
  • 317887 04/17/10 17:48 silwal/R.Mic more changes to the .crl to make scaler run forever
  • 317883 04/17/10 17:21 Roman / Ahme Moller scintillators HV corrected.
  • 317882 04/17/10 17:21 Franklin Staring run 2898 Cosmic running during moller
  • 317881 04/17/10 17:07 Roman / Ahme Moller HV
  • 317878 04/17/10 16:44 Abdurahim compton epics reboot
  • 317877 04/17/10 16:24 Abdurahim Compton epics reboot
  • 317875 04/17/10 15:04 Jack VDC mixing system re-started
  • 317867 04/17/10 12:59 jwexler Changed db files
  • 317866 04/17/10 12:58 O.Glamazdin Helicity board parameter change
  • 317865 04/17/10 12:39 O.Glamazdin Moller magnet alignment test results
  • 317864 04/17/10 12:28 O.Glamazdin Moller magnet has been filled
  • 317863 04/17/10 11:15 dalton Temperatures overnight
  • 317859 04/17/10 10:51 dalton Runs changed to 720000 events
  • 317858 04/17/10 10:43 dalton Prompt and Charge
  • 317857 04/17/10 10:31 silwal dithering turned off
  • 317852 04/17/10 10:31 silwal PC IHWP IN/OUT Voltages and PITA slopes from Injector Studies
  • 317847 04/17/10 09:44 R. Michaels re: charge per run
  • 317845 04/17/10 09:24 O.Glamazdin Moller magnet alignment test
  • 317844 04/17/10 09:18 ZAFAR AHMED The target is empty target
  • 317843 04/17/10 09:13 aniol charge collected per run?
  • 317834 04/17/10 07:58 M. Jen 04/17/10 Owl Shift Summary
  • 317833 04/17/10 07:52 efuchey target GUIs screenshots, target stability
  • 317826 04/17/10 06:57 cusanno End-of-Run 2871 lost
  • 317814 04/17/10 04:56 cusanno minor troubles with CODA
  • 317813 04/17/10 04:53 cusanno Half-Wave-Plate OUT starting from run 2867
  • 317807 04/17/10 04:26 cusanno Dithering
  • 317806 04/17/10 04:23 cusanno spot++
  • 317788 04/17/10 02:24 dalton Half Wave Plate HWP change to out. Slug 5
  • 317776 04/17/10 00:19 M. Jen Hall-A General Tool
  • 317774 04/17/10 00:00 aniol cooldown time for collimators
  • Friday

  • 317772 04/16/10 23:59 Aniol swing shift summary
  • 317771 04/16/10 23:49 Aniol MCC takes away beam
  • 317770 04/16/10 23:34 aniol beam/temp strip chart
  • 317762 04/16/10 22:32 Aniol compton ion chambers tripping beam
  • 317761 04/16/10 22:22 Franklin Coda restart
  • 317756 04/16/10 22:03 dalton Temperature equlibrium at 54 uA
  • 317751 04/16/10 21:04 dalton Sample dithering cycle
  • 317748 04/16/10 20:46 Aniol continue to run at 50uA production
  • 317745 04/16/10 20:35 Aniol temp. history, 50uA
  • 317744 04/16/10 20:18 D. Parno Target temp, vacuum behavior
  • 317737 04/16/10 19:25 dalton Going to 50 uA
  • 317736 04/16/10 19:19 dalton Reset Left Q1 at 17:05
  • 317735 04/16/10 19:12 Aniol going to 50uA
  • 317731 04/16/10 19:08 Guy Ron Going to 50uA
  • 317730 04/16/10 19:04 Jack VDC high voltage should remain OFF
  • 317729 04/16/10 19:01 Franklin VDC Gas mixture problem
  • 317728 04/16/10 18:56 Aniol temp. charts
  • 317725 04/16/10 18:44 Aniol dithering turned on
  • 317714 04/16/10 18:00 D. Parno Target Screen Capture
  • 317712 04/16/10 17:55 riordan GEM Analysis Status
  • 317709 04/16/10 17:39 D. Parno Moved to Pb Diamond Target #3
  • 317708 04/16/10 17:36 aniol spot_R
  • 317707 04/16/10 17:33 aniol spot_R
  • 317703 04/16/10 17:24 silwal modified ~/db/misc/Prex.misc to fill appropriate trees in the PAN output rootfiles
  • 317701 04/16/10 17:18 D. Parno Target to Empty
  • 317700 04/16/10 17:12 Aniol target screen, after LQ1 power cycle
  • 317699 04/16/10 17:06 mercado Database change for BCMs
  • 317698 04/16/10 17:02 aniol Left data suspect after 14:20
  • 317697 04/16/10 17:00 aniol power cycle LQ1
  • 317692 04/16/10 16:43 aniol LQ1 readback = 0A ! page tech
  • 317687 04/16/10 16:09 R. Michaels re: slow EPICS
  • 317682 04/16/10 15:50 silwal dithering plots, bpms
  • 317679 04/16/10 15:16 kowalski PREX
  • 317678 04/16/10 15:08 kowalski PREX
  • 317676 04/16/10 15:00 mcnulty Fixed x distribution plot for Rarm alignment
  • 317675 04/16/10 14:57 silwal using 'fillditheronly true' to thin the datastream
  • 317674 04/16/10 14:35 Guy Ron BCM Calibration
  • 317673 04/16/10 14:30 silwal dithering turned on towards the end of run # 2831
  • 317672 04/16/10 14:30 benesch results of Compton tuning attempt this morning
  • 317667 04/16/10 13:34 M Friend Compton rates after tune
  • 317666 04/16/10 13:13 ZAFAR AHMED Target is Pb Diamon #3
  • 317665 04/16/10 13:07 ZAFAR AHMED Slpt++
  • 317664 04/16/10 13:04 Jack (Techs) More thermocouples on the target
  • 317659 04/16/10 12:55 ZAFAR AHMED Target is BeO
  • 317658 04/16/10 12:49 dalton Epics events slow again
  • 317657 04/16/10 12:37 mcnulty Larm A_T det alignment (update)
  • 317656 04/16/10 12:32 dalton Temperature decay
  • 317655 04/16/10 12:19 O.Glamazdin Moller magnet has been filled
  • 317654 04/16/10 12:18 Abdurahim Compton laser and cavity lock has been restored
  • 317653 04/16/10 12:16 mcnulty Access activity
  • 317652 04/16/10 11:21 kowalski Magnets
  • 317649 04/16/10 10:35 ZAFAR AHMED Dithering is off
  • 317642 04/16/10 09:50 ZAFAR AHMED Target is empty target.
  • 317640 04/16/10 09:36 M Friend Compton EPICS reboot
  • 317639 04/16/10 09:34 Guy Ron New Thermocouples
  • 317628 04/16/10 08:02 Jack (Techs) Logging of temperatures in EPICS
  • 317626 04/16/10 07:48 R. Michaels added temps to Epics logger
  • 317621 04/16/10 07:16 R. Michaels temps at 30 uA (again)
  • 317614 04/16/10 06:05 R. Michaels temps at 30 uA
  • 317613 04/16/10 05:55 R. Michaels attempt #N to fix CH scaler
  • 317609 04/16/10 05:36 silwal bcm1 before/after recent .crl modifications
  • 317606 04/16/10 05:27 silwal more .crl changes to accomodate CH 'dirty trick'
  • 317604 04/16/10 05:05 R. Michaels happex3_ts.crl mod
  • 317603 04/16/10 05:01 efuchey Target GUIs screenshots
  • 317600 04/16/10 04:41 R. Michaels temps at 30 uA (update)
  • 317595 04/16/10 04:03 silwal turned dithering on
  • 317594 04/16/10 03:58 Guy Ron Temp Stripcharts
  • 317589 04/16/10 03:09 adaq Stripcharts
  • 317586 04/16/10 03:03 R. Michaels Temp vs time
  • 317577 04/16/10 02:29 Guy Ron Temperature Stripchart
  • 317574 04/16/10 02:19 jwexler Quartz and Lumi HV
  • 317573 04/16/10 02:17 R. Michaels blinding for Pb ?
  • 317572 04/16/10 02:13 R. Michaels 20 uA on lead
  • 317570 04/16/10 01:56 efuchey Target move to Pb-Diamond
  • 317569 04/16/10 01:48 efuchey Target Move
  • 317566 04/16/10 01:45 Guy Ron First 15 min of heat test
  • 317564 04/16/10 01:41 R. Michaels septum turned on.
  • 317560 04/16/10 01:13 R. Michaels re: runs
  • 317559 04/16/10 01:13 efuchey Target move
  • 317558 04/16/10 01:11 R. Michaels spot++ run
  • 317555 04/16/10 01:00 efuchey Target move
  • 317552 04/16/10 00:57 R. Michaels DAQ mod (late entry)
  • 317551 04/16/10 00:55 riordan Beam Position Scan Summary
  • 317550 04/16/10 00:55 jwexler Changed to integrating mode
  • 317549 04/16/10 00:35 A. Camsonne GEM HV off
  • 317548 04/16/10 00:27 riordan Sieve is OUT
  • 317547 04/16/10 00:14 riordan VDCs, Septum, Upper Scintillator off
  • 317546 04/16/10 00:10 efuchey Target GUIs screenshots
  • 1 04/16/10 00:08 aniol bta hangup
  • 317541 04/16/10 00:00 Aniol swing shift summary
  • Thursday

  • 317535 04/15/10 23:44 aniol r27402 has gem data
  • 317531 04/15/10 23:42 A. Camsonne GEM HV on
  • 317486 04/15/10 22:27 riordan VDCs On, Septum On
  • 317485 04/15/10 22:27 A. Shahinyan Target move
  • 317484 04/15/10 22:26 aniol power permit back
  • 317483 04/15/10 22:12 R. Michaels ask for access, turn off septum
  • 317482 04/15/10 22:09 R. Michaels Pb target position scan, finished.
  • 317481 04/15/10 22:08 A. Shahinyan Target move
  • 317450 04/15/10 21:34 R. Michaels Centering on Pb, part 1
  • 317435 04/15/10 21:16 R. Michaels putting in Pb/diamond at few nA
  • 317434 04/15/10 21:14 A. Shahinyan Target move
  • 317433 04/15/10 21:08 benesch better elog for MCZ1H04 calibration
  • 317431 04/15/10 21:00 A. Shahinyan Target move
  • 317430 04/15/10 20:57 R. Michaels calibration of MCZ1H04 for 3 mm box.
  • 317428 04/15/10 20:44 R. Michaels Beam centering on BeO
  • 317427 04/15/10 20:37 A. Shahinyan Target move
  • 317426 04/15/10 20:17 ZAFAR AHMED Target is BeO
  • 317425 04/15/10 20:16 A. Camsonne GEM HV off
  • 317422 04/15/10 20:12 R. Michaels spot size adjustment
  • 317419 04/15/10 20:08 A. Camsonne GEM HV INFN changed
  • 317412 04/15/10 19:58 ZAFAR AHMED Target is empty target.
  • 317409 04/15/10 19:36 jwexler Detector current 'near final' encoder readouts.
  • 317402 04/15/10 19:18 mcnulty Larm quartz alignment (near final)
  • 317401 04/15/10 19:12 mcnulty Rarm quartz alignment (near final)
  • 317399 04/15/10 19:05 mcnulty Quartz signals
  • 317393 04/15/10 18:36 Jack (Techs) Thermocouple on the beampipe - readout channel corrected
  • 317389 04/15/10 18:30 aniol bta working now
  • 317382 04/15/10 18:28 A. Camsonne GEM HV changed
  • 317381 04/15/10 18:24 A. Shahinyan Target move
  • 317380 04/15/10 18:21 mcnulty spot++
  • 317379 04/15/10 18:20 jwexler BPM values.
  • 317377 04/15/10 18:17 aniol asked for beam permit
  • 317376 04/15/10 18:17 A. Shahinyan Target move
  • 317375 04/15/10 18:16 aniol turned on septum
  • 317374 04/15/10 18:03 aniol septum turned off
  • 317373 04/15/10 18:02 aniol controlled acces
  • 317372 04/15/10 17:59 aniol controlled access
  • 317366 04/15/10 17:39 aniol septum ok
  • 317358 04/15/10 17:28 A. Shahinyan Target move
  • 317357 04/15/10 17:24 aniol septum does not come to 729A
  • 317356 04/15/10 17:20 mcnulty Septum on (but only at 725A)
  • 317355 04/15/10 17:18 mcnulty spot++
  • 317352 04/15/10 17:10 jwexler HVs on
  • 317351 04/15/10 17:05 mcnulty Sieves taken out during access
  • 317350 04/15/10 17:04 mcnulty Access Activity
  • 317349 04/15/10 16:56 A. Camsonne INFN GEMs to 4100 V
  • 317348 04/15/10 16:54 aniol ask for beam permit
  • 317347 04/15/10 16:48 Folts collimator shim
  • 317346 04/15/10 16:45 Aniol turned off septum
  • 317345 04/15/10 16:44 O.Glamazdin Moller magnet has been filled
  • 317344 04/15/10 16:39 A. Shahinyan Target Status
  • 317343 04/15/10 16:36 adaq bta not working
  • 317342 04/15/10 16:35 A.Shahinyan Target Status
  • 317341 04/15/10 16:33 A. Camsonne GEM powered on
  • 317340 04/15/10 16:30 A. Camsonne GEM CAEN HV module power cycled and moved to another NIM slot
  • 317338 04/15/10 16:01 Holmstrom Day Shift Summary
  • 317337 04/15/10 15:42 mercado LED pulse on right arm detector
  • 317336 04/15/10 14:59 D. Parno Target Vacuum Fluctuations (Normal)
  • 317322 04/15/10 13:17 Guy Ron Collimator Temperature
  • 317321 04/15/10 12:48 A. Camsonne GEM HV in weird state
  • 317320 04/15/10 12:40 D. Parno Moved to empty target
  • 317319 04/15/10 12:39 Holmstrom Septum back up
  • 317318 04/15/10 12:22 silwal first look at dithering, dets and slope/position at target
  • 317317 04/15/10 11:49 silwal FFB instabilities??
  • 317316 04/15/10 11:42 mcnulty Raster size correction II
  • 317315 04/15/10 11:02 jwexler Switched to counting mode
  • 317313 04/15/10 10:40 silwal Seive IN
  • 317304 04/15/10 10:20 Holmstrom Turn off Septum Magnet
  • 317303 04/15/10 10:13 silwal CH scaler dead/not working
  • 317296 04/15/10 09:07 D. Parno Restarted Target GUI
  • 317294 04/15/10 08:34 silwal first look at dithering
  • 317292 04/15/10 08:27 silwal RHWP back @ 2350, and feedback on
  • 317290 04/15/10 08:22 D. Parno Target tools
  • 317288 04/15/10 08:03 silwal new 50K 50uA thick Carbon Quartz voltages
  • 317285 04/15/10 07:57 silwal RHWP @ 2900 for big charge asym run, feedback turned off
  • 317284 04/15/10 07:55 M. Jen 4/15/10 Owl Shift Summery
  • 317281 04/15/10 07:37 silwal 50uA Thick Carbon run summary/dithering checkout
  • 317275 04/15/10 06:48 M. Jen turn the septum on; change Lumi HV for Septum on w/ 50uA on thick Carbon target
  • 317266 04/15/10 05:35 A. Camsonne Compton GSO HV Off
  • 317265 04/15/10 05:34 A. Camsonne Compton GSO HV on - Background check
  • 317260 04/15/10 05:18 M. Jen Lumi HV changed for septum off w/ 50uA on thick Carbon target
  • 317259 04/15/10 05:12 M. Jen spot++ w/ 5uA beam, septum off
  • 317255 04/15/10 05:09 A. Camsonne GEM configuration is slow
  • 317254 04/15/10 05:05 silwal 50K @ 50uA on thick Carbon Quartz HVs
  • 317253 04/15/10 05:03 silwal 90-100K @ 50uA on thick Carbon Quartz HVs
  • 317252 04/15/10 04:59 efuchey target move
  • 317251 04/15/10 04:58 M. Jen turn the septum off for the next phase
  • 317250 04/15/10 04:58 silwal Beam Position Scan, +-2mm in x/y on BPM4B, quick check before putting 50uA on Thick Carbon
  • 317249 04/15/10 04:55 silwal S0 turned off on both arms
  • 317248 04/15/10 04:52 M. Jen 5th BPM w/ raster size 8x8
  • 317243 04/15/10 04:46 M. Jen 4th BPM w/ raster size 8x8
  • 317237 04/15/10 04:35 M. Jen 3rd BPM set, raster size 8x8
  • 317232 04/15/10 04:27 M. Jen 2nd BPM w/ raster size 8x8
  • 317231 04/15/10 04:27 silwal 50% deadtime @ 300 Hz, left arm spec daq
  • 317227 04/15/10 04:22 M. Jen 1st BPM w/ raster size 8x8
  • 317220 04/15/10 04:09 silwal S0 turned on, both arms
  • 317219 04/15/10 04:02 M. Jen turn the septum on for dithering test with different BPM conditions
  • 317218 04/15/10 04:00 efuchey target move
  • 317217 04/15/10 03:48 M. Jen integration mode: 5uA, raster on, septum off
  • 317214 04/15/10 03:39 M. Jen turn the Septum off before asking 5uA beam back
  • 317213 04/15/10 03:38 mcnulty Target theta vs detector y at 1.8m
  • 317212 04/15/10 03:36 efuchey target move
  • 317211 04/15/10 03:33 mcnulty Target Angles
  • 317208 04/15/10 03:30 ZAFAR AHMED Cavity Pregains and Phases
  • 317207 04/15/10 03:30 mcnulty xy focal plane dists for both arms
  • 317204 04/15/10 03:26 mcnulty Larm alignment (not complete)
  • 317199 04/15/10 03:12 mcnulty Raster size correction
  • 317198 04/15/10 03:11 A. Camsonne Current GEM database and mapping
  • 317197 04/15/10 03:09 A. Camsonne Current GEM database and mapping
  • 317196 04/15/10 03:08 silwal spot++, 3x4mm, run # 27359
  • 317195 04/15/10 03:07 A. Camsonne Current GEM database and mapping
  • 317194 04/15/10 03:06 mcnulty Rarm alignment (not complete)
  • 317193 04/15/10 03:06 M. Jen Hall-A General Tools
  • 317192 04/15/10 03:06 jwexler HV map for integrating mode, thick Pb and C running.
  • 317190 04/15/10 03:03 A. Camsonne GEM run list
  • 317188 04/15/10 02:59 efuchey target move
  • 317187 04/15/10 02:53 A. Camsonne GEM HV off
  • 317186 04/15/10 02:49 mcnulty Quartz signals
  • 317185 04/15/10 02:44 jwexler Late Entry - VDC, s0, upper scint. off.
  • 317178 04/15/10 02:05 mcnulty Activity during earlier access
  • 317176 04/15/10 01:53 jwexler Det. positions and encoder notes
  • 317161 04/15/10 00:54 mcnulty spot++ raster size check
  • 317159 04/15/10 00:48 fuchey tqrget move
  • 317158 04/15/10 00:47 M. Jen Hall-A General Tool
  • 317153 04/15/10 00:33 efuchey Target move
  • 317149 04/15/10 00:22 efuchey Screenshots: NewDesign and Cryostat
  • Wednesday

  • 317147 04/14/10 23:50 Carlos Swing shift summary
  • 317146 04/14/10 23:20 A . Camsonne GEM HV lowered
  • 317145 04/14/10 23:03 A. Camsonne GEM HV unplugged and replugged USB cable to laptop
  • 317144 04/14/10 22:57 A. Camsonne GEM HV ( slightly more readable )
  • 317116 04/14/10 20:59 jwexler Detectors moved (again)
  • 317111 04/14/10 20:47 jwexler Main det. moved
  • 317107 04/14/10 20:43 benesch spot++ at target suggests different ratio for x,y raster
  • 317103 04/14/10 20:18 jwexler s0, T5 and quartz turned on
  • 317102 04/14/10 20:10 jwexler VDCs turned on
  • 317100 04/14/10 19:51 Carlos Raster & spot++ confusion
  • 317098 04/14/10 19:48 Apar spot++ after adjustment
  • 317095 04/14/10 19:44 Apar spot++ result
  • 317090 04/14/10 19:34 Apar Spot++ result
  • 317088 04/14/10 19:22 adaq Harp scan result for HA1H04B
  • 317087 04/14/10 19:21 adaq Harp scan results for HA1H04A
  • 317079 04/14/10 16:41 A. Camsonne GEM cabling / mapping
  • 317075 04/14/10 16:28 Kiadtisak GEM's HV cables
  • 317073 04/14/10 16:20 aniol target is cold, screen shots
  • 317065 04/14/10 15:58 O.Glamazdin Moller magnet has been filled
  • 317061 04/14/10 15:24 jpchen Alarm limits set.
  • 317060 04/14/10 15:21 jpchen target crystat screen
  • 317059 04/14/10 15:19 jpchen Target cool-down completed
  • 317056 04/14/10 15:06 Jack (Techs) Tungsten o-ring shield
  • 317051 04/14/10 13:57 folts checklist/tech
  • 317034 04/14/10 12:16 jwexler Lower Left Detector Repairs
  • 317032 04/14/10 11:48 R. Michaels scaler fixes; added 2nd clock
  • 317030 04/14/10 11:14 R. Michaels plugged in T8 on R-HRS
  • 317025 04/14/10 10:48 folts Magnets Ramped down
  • 317022 04/14/10 10:34 LeRose Radiation Monitoring
  • 317013 04/14/10 10:21 LeRose Radiation Monitoring
  • 317005 04/14/10 00:49 M Friend raster off
  • Tuesday

  • 317004 04/13/10 23:39 jwexler Left arm quartz detectors turned off
  • 316993 04/13/10 18:10 jwexler Quartz dets. back to HRS DAQ
  • 316992 04/13/10 17:13 Bob/Rupesh investigation of slow E tree data
  • 316987 04/13/10 16:33 A. Camsonne GEM cosmics run 1095
  • 316982 04/13/10 15:10 O.Glamazdin Moller magnet has been filled
  • 316951 04/13/10 11:20 mcnulty switched quartz signals to HRS DAQ
  • 316944 04/13/10 09:22 R. Michaels hypothesis for why O-ring failed
  • Monday

  • 316930 04/12/10 16:27 R. Michaels about O-rings
  • 316929 04/12/10 15:38 O.Glamazdin Moller magnet has been filled
  • 316926 04/12/10 14:58 jwexler Linearity signal before cable swap
  • 316920 04/12/10 14:19 mercado Swapped LEDs. Now get ribbon signals! Added figures..
  • 316919 04/12/10 14:14 mercado Swapped LEDs. Now get ribbon signals!
  • 316917 04/12/10 13:44 mercado Lumi Pedestals - Light Leaks?
  • Sunday

  • 316872 04/11/10 23:35 Abdurahim Searching for optimized DOCP state at the CIP
  • 316865 04/11/10 13:21 O.Glamazdin Moller magnet has been filled
  • Saturday

  • 316820 04/10/10 20:39 Abdurahim Green Compton Cavity Power is 2.5 kW !
  • 316813 04/10/10 16:13 mercado Detector vs BCM fit - pedestal error?
  • 316808 04/10/10 11:41 O.Glamazdin Moller magnet has been filled
  • Friday

  • 316775 04/09/10 15:29 jwexler Mapping between the root files and the BPM patch panel.
  • 316750 04/09/10 07:36 R. Subedi Shift cancelled until Sunday Day-shift
  • Thursday

  • 316743 04/08/10 22:29 O.Glamazdin Moller magnet has been filled
  • 316736 04/08/10 17:42 mercado Summary of what happened during Optics Run 2642-2644
  • 316735 04/08/10 17:38 R. Subedi RHRS evtypebits working
  • 316734 04/08/10 16:34 M Friend Compton Saclay DAQ/CompSimple remote control
  • 316731 04/08/10 14:17 jwexler ADC & BSY BPM patch panel mapping
  • 316729 04/08/10 11:36 dalton Carbon optics data
  • 316718 04/08/10 10:48 R. Subedi Target chamber open: 4th carbon foil broken off
  • 316705 04/08/10 03:21 A. Camsonne Late entry : GEM taken out from Left ARM
  • Wednesday

  • 316698 04/07/10 22:51 R. Michaels upcoming meetings for PREX
  • 316697 04/07/10 22:40 R. Michaels Experiment status and plans
  • 316690 04/07/10 16:27 Kiadtisak Sieve Slit Plots (Tune B)
  • 316689 04/07/10 15:38 Fansler/Smit Moeller
  • 316688 04/07/10 12:28 silwal some injector bpms dropped from the .db file, but still in the raw data file
  • 316680 04/07/10 08:55 R. Subedi Day shift canceled, swing on standby
  • 316679 04/07/10 06:49 R. Subedi A copy from Elog on the early AM events yesterday
  • Tuesday

  • 316678 04/06/10 19:03 riordan LQ1 liquid level dropped with beam
  • 316677 04/06/10 16:36 O.Glamazdin Moller magnet influence on the MBC1H01V corrector
  • 316676 04/06/10 16:27 R. Subedi Shift cancellation and standby
  • 316674 04/06/10 16:03 Fatiha Day Shift Summary
  • 316673 04/06/10 16:00 jpchen only 5 uA on lead target ealier
  • 316672 04/06/10 15:59 Fatiha Lecroy HV alarms
  • 316671 04/06/10 15:04 Fatiha Restricted access
  • 316670 04/06/10 15:02 Fatiha Hall still in controlled access
  • 316669 04/06/10 14:32 Fatiha Hall is going to restricted access
  • 316668 04/06/10 13:15 ZAFAR AHME Target GUI.
  • 316667 04/06/10 12:49 Fatiha Hall A General tools
  • 316666 04/06/10 12:42 Fatiha 1/2 Day Shift Summary
  • 316665 04/06/10 12:42 dalton Shift Summary Owl
  • 316664 04/06/10 12:39 pking Beam current and radiation monitor levels from archiver; comparison to target motion logs
  • 316663 04/06/10 12:33 O. Glamazdin Moller magnet is full with LHe
  • 316662 04/06/10 12:32 jpchen Chamber vacuum bad, suspect leak.
  • 316661 04/06/10 12:09 Fatiha More people in the hall
  • 316660 04/06/10 11:52 Fatiha A couple of people in the hall
  • 316659 04/06/10 11:22 jpchen scattering chamber vacuum still bad
  • 316658 04/06/10 10:29 Fatiha Askd for a controlled access
  • 316657 04/06/10 10:20 R. Michaels no thin lead in beam, please !
  • 316656 04/06/10 09:50 jpchen Moved to BeO
  • 316655 04/06/10 09:44 jpchen beam on thick 12C target caused the vacuum bad
  • 316654 04/06/10 09:37 Moffit bpm saturation cuts in multirun analysis
  • 316647 04/06/10 08:43 ZAFAR AHMED Target temp is increasing
  • 316646 04/06/10 08:37 dalton Target vacuum stops beam
  • 316644 04/06/10 08:31 ZAFARA AHMED Vacuum pressure alarm
  • 316643 04/06/10 08:24 D. Armstrong late follow up on spot++ on Pb target
  • 316642 04/06/10 08:21 ZAFAR AHMED Target alarm handler red
  • 316640 04/06/10 08:03 mercado Detector HV Adjust for Carbon
  • 316639 04/06/10 08:01 A. Giusa Vacuum pressure problem
  • 316636 04/06/10 07:20 A. Giusa Target moved to thick Carbon
  • 316634 04/06/10 07:10 dalton Quartz HV
  • 316633 04/06/10 07:08 dalton Quartz HV
  • 316632 04/06/10 06:43 silwal "old" Pb#3 runs @ 50uA
  • 316631 04/06/10 06:42 silwal "old" Pb#3 runs @ 50uA
  • 316629 04/06/10 06:42 dalton Changing to HWP IN
  • 316624 04/06/10 06:06 dalton RHWP to 2900 and charge feedback off
  • 316618 04/06/10 05:00 dalton Left Dipole lock
  • 316608 04/06/10 04:35 A. Giusa Vacuum pressure Ok
  • 316603 04/06/10 03:44 A.Giusa Vacuum pressure
  • 316602 04/06/10 03:42 dalton Halog run start bug fixed
  • 316601 04/06/10 03:39 A. Camsonne Compton GSO HV off
  • 316598 04/06/10 03:22 A. Camsonne Compton rates too high in transverse setting
  • 316596 04/06/10 03:12 A. Giusa Grab of target screen
  • 316593 04/06/10 02:49 dalton spot++ on Pb diamond #3
  • 316592 04/06/10 02:49 A. Giusa Target moved to Pb Diamond #3
  • 316590 04/06/10 02:48 dalton thin lead
  • 316586 04/06/10 02:38 dalton spot++
  • 316577 04/06/10 02:23 A. Giusa Vacuum going Ok
  • 316576 04/06/10 02:19 silwal curmon changed from bcm3 to bcm1
  • 316574 04/06/10 02:08 silwal PITA scan, IHWP IN, RHWP @ 2350, run # 2644
  • 316572 04/06/10 01:47 Apar PITA scan, IHWP out, RHWP @ 2350, slope ~ 8 ppm/V
  • 316571 04/06/10 01:39 dalton Left Q1 quench (tech)
  • 316570 04/06/10 01:17 A. Giusa Minor Alarm on Scattering Vacuum
  • 316568 04/06/10 01:00 R. Michaels re: spot at 5 nA
  • 316560 04/06/10 00:48 A. Giusa Yellow alarm on scattering channel vaccum
  • 316559 04/06/10 00:45 A. Giusa Target on optics
  • 316554 04/06/10 00:25 D. Armstrong Shift Summary - Swing shift
  • 316552 04/06/10 00:21 Armstrong spot++ on Pb target
  • 316551 04/06/10 00:17 D. Armstrong 5nA on Pb target #3
  • 316549 04/06/10 00:13 R. Michaels re: rate at low current
  • 316548 04/06/10 00:04 D. Armstrong 5 nA (?) on BeO
  • Monday

  • 316546 04/05/10 23:41 D. Armstrong go to BeO
  • 316545 04/05/10 23:38 A. Camsonne Compton GSO HV on at 1900 V
  • 316544 04/05/10 23:36 Armstrong spot++
  • 316538 04/05/10 23:27 Sarah Hall A General Tools Screen Snapshot
  • 316537 04/05/10 23:22 D. Armstrong 5 uA in Hall
  • 316536 04/05/10 23:16 D. Armstrong Start up Pb target recover procedure
  • 316535 04/05/10 22:35 D. Armstrong Beam soon
  • 316534 04/05/10 22:31 O. Glamazdin Moller results for Spin Dance
  • 316533 04/05/10 22:16 romanip The BCM output is diconnected from Moller DAQ's
  • 316531 04/05/10 21:03 D. Armstrong Spin dance completed. Ask for controlled access
  • 316530 04/05/10 20:22 D. Armstrong Status update
  • 316529 04/05/10 19:08 E.Chudakov BCM x10 linearity
  • 316528 04/05/10 19:01 D. Armstrong Status update
  • 316520 04/05/10 16:01 Fatiha Day shift summary
  • 316519 04/05/10 15:45 A. Giusa grab of target screen
  • 316518 04/05/10 14:49 Sirish Compton beam polarimetry
  • 316517 04/05/10 14:45 LeRose Alternate Tunes
  • 316516 04/05/10 14:45 silwal this is bcm9
  • 316515 04/05/10 14:44 A. Camsonne Compton eventSelection updated
  • 316514 04/05/10 14:14 romanip The BCM output for the Moller DAQ's
  • 316513 04/05/10 13:51 O.Glamazdin Moller magnet ioc was rebooted
  • 316512 04/05/10 13:43 A. Camsonne Compton GSO HV off
  • 316511 04/05/10 13:13 Jixie Hall A General Tools
  • 316510 04/05/10 13:09 Jixie test
  • 316509 04/05/10 13:07 romanip BCM output for Moller DAQ's
  • 316508 04/05/10 12:45 O.Glamazdin Moller magnet field
  • 316507 04/05/10 12:31 Fatiha Spin dance going on
  • 316504 04/05/10 11:39 Fatiha No beam
  • 316503 04/05/10 11:23 Fatiha No beam
  • 316500 04/05/10 10:25 silwal feedback switched from bcm2 to bcm1
  • 316499 04/05/10 10:20 Fatiha End of controlled access
  • 316497 04/05/10 09:38 R. Michaels cleanup of /adaql*/work*
  • 316493 04/05/10 09:24 Fatiha Hall going to escorted access
  • 316482 04/05/10 08:32 Fatiha All magnets down
  • 316480 04/05/10 08:26 pking Owl shift summary
  • 316477 04/05/10 08:15 pking Left Q1 liquid level now shows zero (Techs)
  • 316476 04/05/10 08:12 pking Left Q1 trip: Trying reset proceedure from Prex op manual; didn't reset.
  • 316474 04/05/10 08:06 pking Left Q1 has dropped to zero. (Techs)
  • 316467 04/05/10 06:55 pking Are the BPM wire signals normalized by integration time?
  • 316466 04/05/10 06:46 pking Increased beam current to 60 uA on MPS BCM
  • 316465 04/05/10 06:43 M Friend Compton photon beam position scans
  • 316458 04/05/10 05:46 pking The disk /adaql3/work1 is full; prompt analyses are failing
  • 316455 04/05/10 05:36 M Friend Compton striptool snapshot
  • 316450 04/05/10 04:34 M Friend Cavity stopped locking again
  • 316443 04/05/10 03:58 M Friend Compton cavity QWP motion reboot
  • 316438 04/05/10 03:03 pking IOC se8 and se10 back in autogain
  • 316436 04/05/10 03:01 pking FFB and iocse9 settings after BCM/BPM calibration
  • 316434 04/05/10 02:58 silwal/pking BCM/BPM calibration run # 2617
  • 316430 04/05/10 02:23 Abdurahim Green Cavity Polarization Study
  • 316427 04/05/10 02:04 pking IOC se9 with low RF gain
  • 316426 04/05/10 01:59 pking IOC se8 and se10 set to "gains off"
  • 316422 04/05/10 01:44 silwal bcm1-2 swapped
  • 316418 04/05/10 01:08 pking BPM12 wire signals are very low
  • Sunday

  • 316402 04/04/10 23:58 D. Armstrong Shift Summary - Swing shift
  • 316401 04/04/10 23:54 M Friend Strange looking BPM12x
  • 316396 04/04/10 23:32 A. Camsonne Compton electron detector currents position -2 mm 50 uA
  • 316393 04/04/10 23:20 A. Camsonne Compton GSO detector moved out
  • 316392 04/04/10 23:19 A. Camsonne GEM HV off
  • 316390 04/04/10 23:09 M Friend Compton photon detector PMT replaced
  • 316387 04/04/10 22:52 ZAFAR AHMED Moved to empty target.
  • 316386 04/04/10 22:50 D. Armstrong Controlled Access
  • 316385 04/04/10 22:47 silwal beam modulation gui obsolete
  • 316383 04/04/10 22:04 D. Armstrong Run List for horizontal position scan
  • 316345 04/04/10 20:55 A. Camsonne Compton electron detector in position -2 mm
  • 316313 04/04/10 20:06 D. Armstrong Run list from y-position scan with Une Y
  • 316254 04/04/10 18:47 A. Camsonne Compton electron detector back in garage position
  • 316251 04/04/10 18:45 A. Camsonne Electron detector spectrum
  • 316250 04/04/10 18:44 ZAFAR AHMED Target is supre thin carbon.
  • 316237 04/04/10 18:30 Armstrong Tune "Y" settings
  • 316235 04/04/10 18:28 D. Armstrong previous runs were tune B, not tune Y
  • 316229 04/04/10 18:17 A. Camsonne Current and rates electron detector position -1 mm
  • 316228 04/04/10 18:09 A. Camsonne Compton electron detector to position 0 mm
  • 316218 04/04/10 17:51 A. Camsonne Compton electron detector current at 7 mm from the beam
  • 316217 04/04/10 17:51 D. Armstrong Quartz HV settings
  • 316207 04/04/10 17:33 Armstrong setting up for optics.
  • 316206 04/04/10 17:32 A. Camsonne Requested to mask electron detector FSD for current checks
  • 316197 04/04/10 17:22 A. Camsonne rebooted iocha22
  • 316196 04/04/10 17:09 A. Camsonne GEM HV on for data taking during optics data
  • 316195 04/04/10 16:57 D. Armstrong move to thin Carbon target
  • 316188 04/04/10 16:50 D. Armstrong Access to insert sieves
  • 316187 04/04/10 16:45 A. Camsonne Compton electron detector field probe on field error
  • 316184 04/04/10 15:46 M. Jen 04/04/10 Day Shift Summary
  • 316183 04/04/10 15:28 A. Giusa Target moved to empty position
  • 316182 04/04/10 15:27 M. Jen turn the septum on
  • 316180 04/04/10 15:17 silwal RHWP @ 2350, dithering off, septum on, feedback on
  • 316179 04/04/10 15:14 silwal RHWP @ 2900, dithering off, septum on, feedback off
  • 316178 04/04/10 15:11 silwal lumi/beam widths check; day shift semi-summary
  • 316177 04/04/10 14:57 mercado A_T Scan Results
  • 316175 04/04/10 14:47 M. Jen turn the septum off
  • 316172 04/04/10 14:30 silwal septum off
  • 316171 04/04/10 14:30 silwal septum on, lumi ~ 150K
  • 316167 04/04/10 14:00 A. Giusa target screen grab
  • 316161 04/04/10 12:46 silwal RHWP @ 2900
  • 316154 04/04/10 11:57 M. Jen Beam Lumi distributions with time
  • 316153 04/04/10 11:56 M. Jen adjust Beam Lumi HV on thick Carbon Target w/ 50uA
  • 316150 04/04/10 11:22 M. Jen (4)Quartz HV signal plots for det1/2/3/4/A_T
  • 316149 04/04/10 11:21 M. Jen (3)Quartz Signal in det1/2/3/4/A_T
  • 316148 04/04/10 11:18 M. Jen (2)Quartz HV setting; HV signal plots for det#1/2/3/4/A_T
  • 316147 04/04/10 11:16 M. Jen (1) Quartz HV setting; HV signal plots for det1/2/3/4/A_T
  • 316145 04/04/10 11:00 M. Jen turn Septum on (beam off)
  • 316140 04/04/10 10:22 M. Jen Lumi HV after adjustment
  • 316134 04/04/10 09:49 M. Jen L-arm Q1 quench back to normal now
  • 316130 04/04/10 09:26 jwexler A_T scan summary
  • 316116 04/04/10 08:30 M. Jen L-Q1 quench
  • 316110 04/04/10 08:06 bogdanw shift summary
  • 316107 04/04/10 07:55 adaq Q1 trip He level
  • 316065 04/04/10 03:48 Apar raster on Pb#3
  • 316063 04/04/10 03:40 jwexler Helicity board frequency changed again.
  • 316060 04/04/10 03:06 jwexler Target: BeO in.
  • 316058 04/04/10 02:51 jwexler Helicity board frequency changed to 240Hz
  • 316057 04/04/10 02:42 Apar raster check
  • 316053 04/04/10 02:17 A. Camsonne GEM HV ramped down
  • 316052 04/04/10 02:15 ZAFAR AHMED Changed the quartz signals from counting to integrating mode.
  • 316051 04/04/10 01:51 A. Camsonne Magnetic Field probe for Electron Detector plugged back
  • 316050 04/04/10 01:36 dalton A_t hole on lead
  • 316045 04/04/10 01:22 jwexler A_T detectors moved
  • 316044 04/04/10 01:09 A. Camsonne GEM HV back on at 3700 V
  • 316043 04/04/10 01:08 A. Camsonne Compton electron detector motion test
  • 316042 04/04/10 00:59 jwexler Magnet cycling (tune Y->X->B)
  • 316023 04/04/10 00:29 A. Camsonne GEM HV off
  • 316022 04/04/10 00:28 jwexler Main detectors unchanged from April 2nd.
  • 316021 04/04/10 00:28 A. Camsonne GEM pedestal file updated to run 27274
  • 316009 04/04/10 00:06 D. Armstrong Shift Summary - Swing shift
  • 316008 04/04/10 00:06 dalton Run-plan for tonight
  • 316007 04/04/10 00:02 A. Shahinyan Target position
  • Saturday

  • 316006 04/03/10 23:47 jwexler Quartz, s0 and upper scintillator turned on.
  • 316005 04/03/10 23:30 ZAFAR AHMED Changed of connections Integrating to counting
  • 316004 04/03/10 22:33 O. Glamazdin Moller target commissioning results
  • 316003 04/03/10 22:28 D. Armstrong Moller commissioning studies complete for now
  • 316002 04/03/10 22:27 A. Shahinyan Target status
  • 316001 04/03/10 22:08 mcnulty Comparison between A_T plug blocked and unblocked for tune A, sieve out, 9mg C
  • 316000 04/03/10 21:12 D. Armstrong Moller magnet helium level now 70%
  • 315999 04/03/10 20:52 D. Armstrong Another controlled access
  • 315998 04/03/10 20:29 riordan Right arm is good too
  • 315997 04/03/10 20:19 riordan Tune B Left Arm Sieve very clear with proper VDC position offsets
  • 315996 04/03/10 19:54 D. Armstrong More on Moller solenoid He filling
  • 315995 04/03/10 19:42 D. Armstrong Access over, but Moller magnet dewar not full
  • 315994 04/03/10 19:31 Armstrong New Spectrometer settings
  • 315993 04/03/10 19:24 A. Camsonne GEM pedestal run 27274
  • 315991 04/03/10 19:22 riordan VDC databases updated with focal plane offsets
  • 315989 04/03/10 19:19 A. Camsonne GEM HV on
  • 315988 04/03/10 19:02 A. Camsonne Compton electron detector bogus current readings
  • 315987 04/03/10 18:48 D. Armstrong Controlled access - refill Moller dewar
  • 315986 04/03/10 18:03 Jack (Techs) Magnet problems this morning
  • 315985 04/03/10 17:53 Jack (Techs) Shortcomings of the new pagers
  • 315975 04/03/10 16:46 A. Camsonne Scintillator HV ON for GEM check
  • 315971 04/03/10 15:50 M. Jen 04/03/10 Day Shift Summary
  • 315970 04/03/10 14:55 silwal bpm 8/10/12 widths, run # 2529, multiplet trees
  • 315969 04/03/10 14:53 silwal injector/target bpm widths, run#2529,multiplet tree
  • 315968 04/03/10 14:44 A. Camsonne CRAMS left arm GEM pedestal updated
  • 315967 04/03/10 14:02 A. Camsonne Other pedestal run
  • 315966 04/03/10 13:56 A. Camsonne Other pedestal run
  • 315965 04/03/10 13:49 A. Camsonne Compton GSO HV off
  • 315964 04/03/10 13:10 M Friend Compton FADC DAQ now running at 240Hz
  • 315963 04/03/10 11:33 dalton New pagers confusing (Tech)
  • 315962 04/03/10 11:22 O.Glamazdin The Moller magnet refill
  • 315961 04/03/10 11:10 mercado A_t Detector Scan Results
  • 315957 04/03/10 10:23 mercado BCM pedestal fix
  • 315955 04/03/10 10:12 Antonio Gius Target on empty position
  • 315954 04/03/10 09:40 folts/tech sieve alignment to handle
  • 315953 04/03/10 09:20 jwexler Current A-T hole detector positions
  • 315937 04/03/10 08:34 Apar Helicity @ 240Hz
  • 315935 04/03/10 08:22 O. Glamazdin Moller BCM signal was disconnected
  • 315934 04/03/10 08:17 Antonio Gius Target on Pb Diamond #3
  • 315930 04/03/10 07:58 jwexler Shift Summary - Owl shift 04/03/2010
  • 315928 04/03/10 07:40 A. Camsonne Raster check run 27270
  • 315926 04/03/10 07:32 A. Camsonne Gem pedestal run 27269
  • 315923 04/03/10 07:22 jwexler Controlled access summary.
  • 315922 04/03/10 07:21 A. Camsonne UVA GEM back in place - one gassiplex board replaced
  • 315921 04/03/10 07:15 Ahmed / Rupe swithced the quartz signals from counting mode to integrating mode
  • 315919 04/03/10 06:20 jwexler VDC, s0 and scint. off, quartz HV turned down.
  • 315913 04/03/10 05:54 jwexler Main detectors returned to starting positions
  • 315912 04/03/10 05:53 jwexler Main detectors returned to starting positions
  • 315903 04/03/10 05:32 jwexler Motors moved again
  • 315898 04/03/10 05:24 jwexler Lumis off
  • 315893 04/03/10 05:18 silwal septum fault??/tech
  • 315891 04/03/10 05:12 R. Michaels Sieves removed at 3:15 am (LATE entry)
  • 315885 04/03/10 05:02 jwexler Detectors moved.
  • 315876 04/03/10 04:39 R. Michaels Lead target in beam
  • 315873 04/03/10 04:32 jwexler HV for s0, upper and lower scint. and quartz on.
  • 315850 04/03/10 04:00 R. Michaels attenuator and slit
  • 315849 04/03/10 03:52 jwexler Position and raster for thick lead
  • 315847 04/03/10 03:36 R. Michaels tweak raster to 18 x 8 MCC units
  • 315845 04/03/10 03:32 jwexler Magnet changes during Bob's optics plan
  • 315843 04/03/10 03:31 R. Michaels spot++ check -- slightly big
  • 315841 04/03/10 02:59 R. Michaels Optics runs on Ta for tune B
  • 315817 04/03/10 00:56 A. Camsonne Compton asymmetry run 22437
  • 315806 04/03/10 00:42 M Friend Compton rates with 2cm collimator
  • Friday

  • 315804 04/02/10 23:59 D. Armstrong Shit Summary - Swing Shift
  • 315803 04/02/10 23:49 A. Camsonne GEM check - one group bad needs to be replaced
  • 315802 04/02/10 23:46 Jin Huang Target Screen Shot
  • 315801 04/02/10 23:28 R. Michaels optics run plan from Nilanga
  • 315793 04/02/10 22:40 adaq D. Armstrong
  • 315792 04/02/10 22:33 R. Michaels sieves put in
  • 315791 04/02/10 22:27 dalton Runplan tonight
  • 315790 04/02/10 22:11 riordan Right Arm y Shoulder
  • 315787 04/02/10 21:22 M Friend Compton background rates
  • 315786 04/02/10 21:21 D. Armstrong reboot IOCs
  • 315785 04/02/10 21:19 M Friend Counting house BCMs
  • 315784 04/02/10 20:52 D. Armstrong ask for 100 µA for more Compton study
  • 315783 04/02/10 20:31 D. Armstrong Access completed
  • 315782 04/02/10 20:20 D. Armstrong Access completed
  • 315781 04/02/10 19:55 M Friend Compton photon collimator changed
  • 315780 04/02/10 19:17 Jin Huang Update VDC database: T0 update from Seamus
  • 315778 04/02/10 18:48 silwal BCM/BPM calibrated
  • 315776 04/02/10 17:18 adaq Green Compton Commissioning
  • 315762 04/02/10 15:59 Saha Shift Summary
  • 315761 04/02/10 14:57 R. Michaels spot for 100uA test (historical)
  • 315760 04/02/10 14:41 paschke Detector footprint, Right arm, after 4/2 alignment. 4x4 raster
  • 315759 04/02/10 14:37 paschke Detector footprint, Right arm, after 4/2 alignment. 3x3 raster
  • 315758 04/02/10 14:31 paschke Detector footprint, Right arm, after 4/2 alignment. Tiny raster
  • 315757 04/02/10 14:22 paschke Detector footprint, left arm, after 4/2 alignment. 4x4 raster
  • 315756 04/02/10 14:20 paschke Detector footprint, left arm, after 4/2 alignment. 3x3 raster
  • 315755 04/02/10 14:16 paschke Detector footprint, left arm, after 4/2 alignment. Small raster.
  • 315754 04/02/10 14:06 R. Michaels recovered March halog
  • 315753 04/02/10 13:57 paschke my notes on spectrometer studies / detector alignment run list, 4/2 Owl
  • 315752 04/02/10 13:39 mercado HAPPEX DAQ Sync Monitor
  • 315744 04/02/10 13:08 adaq test
  • 315738 04/02/10 11:27 M Friend Helicity board setting to 30Hz
  • 315736 04/02/10 11:11 A. Camsonne UVA chamber pedestal width with grounding and signals
  • 315735 04/02/10 11:02 tiefen 9th dipole scanner successful scan
  • 315734 04/02/10 10:25 paschke Detector aligned to thin C elastic peak at end of Owl shift
  • 315721 04/02/10 07:52 ZAFAR AHMED April 2nd Owl Shift Summary
  • 315720 04/02/10 07:36 ZAFAR AHMED Controlled access
  • 315712 04/02/10 06:21 paschke current calibration - run 2524
  • 315704 04/02/10 05:38 ZAFAR AHMED Moved to empty target
  • 315695 04/02/10 05:09 jwexler Right det. moved.
  • 315688 04/02/10 04:40 jwexler Movers move detector position in y again.
  • 315683 04/02/10 04:28 mcnulty y distributions for different projection distances
  • 315682 04/02/10 04:25 jwexler Movers moved both detectors towards the beamline
  • 315675 04/02/10 04:06 jwexler Motors moved detectors 2cm away from target.
  • 315638 04/02/10 02:30 ZAFARA AHMED Target is thin carbon.
  • 315630 04/02/10 01:53 paschke Valve investigation
  • 315629 04/02/10 01:43 mcnulty Another y alignment plot using different projection algorithm
  • 315623 04/02/10 00:57 paschke someone on the way in to look at valve, now.
  • 315622 04/02/10 00:47 mcnulty Here is the y alignment plot (finally)
  • 315621 04/02/10 00:47 mcnulty Here is the y alignment plot
  • 315620 04/02/10 00:43 adaq Partial swing shift summary
  • 315619 04/02/10 00:42 pking CH scalers are working in run 2519 and beyond.
  • 315618 04/02/10 00:41 mcnulty y alignment plot for Left arm
  • 315609 04/02/10 00:16 A. Camsonne GEM HV lowered to 3650 V
  • 315608 04/02/10 00:09 A. Camsonne GEM HV ON
  • Thursday

  • 315601 04/01/10 23:33 A. Camsonne Pedestal file updated for GEM ROC6
  • 315600 04/01/10 23:32 ZAFAR AHMED We moved to empty target
  • 315599 04/01/10 23:08 pking BPM12 was saturated in 2510.
  • 315598 04/01/10 23:01 paschke valve problem on beamline 1H00, no beam
  • 315597 04/01/10 22:52 pking Changed raster to be about 0.5mm by 0.5 mm
  • 315594 04/01/10 22:37 pking Rates for 4 nA on thin carbon
  • 315587 04/01/10 22:23 mcnulty MCC having trouble delivering beam
  • 315586 04/01/10 22:14 pking Rates with no beam
  • 315585 04/01/10 22:10 silwal bmw phase map for each bmw sinewave / CH scalers not working
  • 315584 04/01/10 22:08 Solvignon Target in Thin carbon position
  • 315583 04/01/10 22:00 pking spot++ check from 27212
  • 315580 04/01/10 21:54 mcnulty VDC, s0, and upper scint HV on and quartz HV=-2000V
  • 315579 04/01/10 21:53 mcnulty Access activity: all quartz dets switched to HRS sampling DAQ
  • 315571 04/01/10 21:14 pking Both HRS DAQs are in "NoGem" configuration.
  • 315568 04/01/10 21:05 atrig Trigger-Download
  • 315559 04/01/10 20:37 R. Michaels tune B and "final" septum set
  • 315557 04/01/10 20:26 pking Lots of error messages in ROC6 (happex11)
  • 315556 04/01/10 20:26 riordan Septum Current Adjustment
  • 315553 04/01/10 19:57 pking ROC7 is the GEM crate in RHRS
  • 315547 04/01/10 19:38 pking Tried to cycle hareboot22, outlet 8. ROC7 does not reboot.
  • 315543 04/01/10 19:05 pking RHRS DAQ unable to connect to ROC7
  • 315542 04/01/10 18:53 pking Spot_L for 27207
  • 315539 04/01/10 18:47 pking Spot_L from run 27206
  • 315517 04/01/10 16:03 X. Zheng Day shift summary 4/1
  • 315516 04/01/10 15:13 Jack (Techs) Septum magnet & Moller Target magnet ramp down if there is an access
  • 315513 04/01/10 14:56 mcnulty y distributions from Bob's septum scan done on 3/31 with central hole cut
  • 315512 04/01/10 13:41 X. Zheng Beam study/access update
  • 315505 04/01/10 13:23 mcnulty y distributions from Bob's septum scan done on 3/31
  • 315504 04/01/10 13:14 X. Zheng MCC: Expect beam in ~15 minutes
  • 315501 04/01/10 13:07 mercado Pedestals make sense now.
  • 315498 04/01/10 12:59 mercado DAQ tests.
  • 315497 04/01/10 12:55 A. Camsonne Compton electron detector work
  • 315496 04/01/10 12:54 mercado Recalibrated some CH crate 18-bit ADC pedestals.
  • 315494 04/01/10 12:27 X. Zheng Beam study/access update - another 1/2 or 1 hour
  • 315492 04/01/10 12:21 A. Camsonne Compton electron detector current no beam
  • 315481 04/01/10 11:45 A. Camsonne Compton electron detector noise check
  • 315478 04/01/10 11:27 mercado Current BCM Configuration
  • 315477 04/01/10 11:17 O.Glamazdin Moller FADC aperture chanels swapped
  • 315476 04/01/10 11:05 X. Zheng Beam study/access update - at least another hour
  • 315475 04/01/10 11:03 X. Zheng Hall A Tools
  • 315472 04/01/10 10:19 X. Zheng full survey is finished
  • 315470 04/01/10 09:29 X. Zheng full survey of the hall is underway
  • 315462 04/01/10 09:10 mcnulty Called for controlled access (9:10am)
  • 315454 04/01/10 07:57 jwexler Shift Summary for Owl 04/01/2010
  • 315449 04/01/10 07:14 R. Michaels scans for tune A,B: plots and discussion
  • 315446 04/01/10 06:54 M. Jen BPM/BCM in pseudo-production mode
  • 315442 04/01/10 06:23 jwexler Shift Checklist: Target Screen - 04/01/10
  • 315440 04/01/10 06:21 jwexler Shift checklist: Hall A tools screen - Owl 04/01
  • 315438 04/01/10 06:00 R. Michaels back out from scans
  • 315437 04/01/10 05:57 R. Michaels Q1 scan, tune A
  • 315376 04/01/10 05:28 R. Michaels septum scan, tune A
  • 315331 04/01/10 05:04 jwexler Entered tune A
  • 315330 04/01/10 04:34 R. Michaels scan of Q1, tune B
  • 315279 04/01/10 04:11 jwexler Corrections to run settings for 6488-98, 27153-27163
  • 315272 04/01/10 04:02 R. Michaels scan of septum, tune B
  • 315227 04/01/10 03:40 R. Michaels HV mapping
  • 315226 04/01/10 03:16 paschke / si First step of dithering commissioning
  • 315204 04/01/10 01:25 paschke HALOG MALFUNCTION - March log entries unavailable
  • 315202 04/01/10 01:10 paschke bpm datamap fix removes raster effect
  • 315194 04/01/10 00:17 pking Swing shift summary